### v3

# A max filtering local stability theorem with application to weighted phase retrieval and cryo-EM

Abstract
Given an inner product space $V$ and a group $G$ of linear isometries, max filtering offers a rich class of convex $G$-invariant maps. In this paper, we identify sufficient conditions under which these maps are locally bilipschitz on $R(G)$, the set of orbits with maximal dimension, with respect to the quotient metric on the orbit space $V/G$. Central to our proof is a desingularization theorem, which applies to open, dense neighborhoods around each orbit in $R(G)/G$ and may be of independent interest.
As an application, we provide guarantees for stable weighted phase retrieval. That is, we construct componentwise convex bilipschitz embeddings of weighted complex (resp. quaternionic) projective spaces. These spaces arise as quotients of direct sums of nontrivial unitary irreducible complex (resp. quaternionic) representations of the group of unit complex numbers $S^{1}\cong\operatorname{SO}(2)$ (resp. unit quaternions $S^{3}\cong\operatorname{SU}(2)$).
We also discuss the relevance of such embeddings to a nearest-neighbor problem in single-particle cryogenic electron microscopy (cryo-EM), a leading technique for resolving the spatial structure of biological molecules.
1 Introduction
Machine learning algorithms are often designed for Euclidean data, represented as vectors in an inner product space $V$. For instance, fast randomized nearest neighbor algorithms, such as the one in [14], efficiently approximate nearest neighbors in large Euclidean datasets by avoiding explicit computation of all pairwise distances.
However, many practical data representations in $V$ involve ambiguities arising from an orthogonal symmetry group $G\leq\operatorname{O}(V)$. For example, as discussed in [19, 20], data from cryogenic electron microscopy (cryo-EM) may reside in a finite-dimensional complex vector space $\mathbb{C}^{d}$, subject to an ambiguity induced by a diagonal circle action $S^{1}\to\mathbb{C}^{d\times d}$ defined by $\theta\mapsto\operatorname{diag}\{e^{\mathrm{i}k_{j}\theta}\}_{j=1}^{d}$, where $\{k_{j}\}_{j=1}^{d}$ are arbitrary fixed integers. This example is elaborated in Section 1.2. (Throughout the paper, we view $\mathbb{C}^{d}$ as a real vector space with inner product $\langle z,x\rangle:=\operatorname{Re}(z^{*}x)$. In particular, the aforementioned action is indeed real-orthogonal.)
To address such ambiguities, one represents the data unambiguously as orbits $x]:=G\cdot x$ in the quotient space $V/G$, equipped with the quotient metric
$d([x],[y]):=\inf_{\begin{subarray}{c}p\in[x]\\ q\in[y]\end{subarray}}\|p-q\|.$
(Indeed, this metric is nondegenerate provided the $G$-orbits are topologically closed). To leverage the extensive machinery of Euclidean-based machine learning, it is desirable to embed the orbit space into Euclidean space in a bilipschitz manner. Specifically, we seek a map $f\colon V/G\to\mathbb{R}^{n}$ and constants $\alpha,\beta>0$ such that
$\alpha\cdot d([x],[y])\leq\|f([x])-f([y])\|\leq\beta\cdot d([x],[y])\qquad\forall x,y\in V.$
The bilipschitz requirement ensures that distances in $V/G$ are faithfully preserved, enabling robust transfer of Euclidean algorithms to the orbit space. For example, adapting the $\lambda$-approximate nearest neighbor problem to $V/G$ becomes straightforward when such embeddings are available.
Example 1 (Example 1 in [8]).
Given $\lambda\geq 1$ and data $x_{1}],\dots,[x_{m}]\in V/G$, the $\lambda$-approximate nearest neighbor problem takes as input $x]\in V/G$ and outputs $j\in\{1,\dots,m\}$ such that
$d([x],[x_{j}])\leq\lambda\cdot\min_{1\leq i\leq m}d([x],[x_{i}]).$
Given a map $f\colon V/G\to\mathbb{R}^{n}$ with bilipschitz bounds $\alpha,\beta\in(0,\infty)$ and given a black box algorithm that solves the problem in $\mathbb{R}^{n}$, one may transfer the algorithm to $V/G$ by pulling back through $f$. This results in a solution of the $\frac{\beta}{\alpha}\lambda$-approximate nearest neighbor problem in $V/G$. To see this, first use the black box algorithm to find $j\in\{1,\dots,m\}$ such that
$\|f([x])-f([x_{j}])\|\leq\lambda\cdot\min_{1\leq i\leq m}\|f([x])-f([x_{i}])\|.$
Then
$d([x],[x_{j}])\leq\frac{1}{\alpha}\cdot\|f([x])-f([x_{j}])\|\leq\frac{\lambda}{\alpha}\cdot\min_{i\in I}\|f([x])-f([x_{i}])\|\leq\frac{\beta}{\alpha}\lambda\cdot\min_{i\in I}d([x],[x_{i}]).$
To this end, [9] recently introduced a family of embeddings called max filter banks that enjoy explicit bilipschitz bounds whenever $G$ is finite. Later work improved on those bounds [17, 18].
Definition 2.
Consider any real inner product space $V$ and $G\leq\operatorname{O}(V)$.
  • (a)
    The max filtering map $\langle\hskip-2.5pt\langle\cdot,\cdot\rangle\hskip-2.5pt\rangle\colon V/G\times V/G\to\mathbb{R}$ is defined by
    $\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle:=\sup_{\begin{subarray}{c}p\in[x]\\ q\in[z]\end{subarray}}\langle p,q\rangle.$
  • (b)
    Given templates $z_{1},\ldots,z_{n}\in V$, the corresponding max filter bank $\Phi\colon V/G\to\mathbb{R}^{n}$ is defined by
    $\Phi([x]):=\{\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle\}_{i=1}^{n}.$
(Since $G\leq\operatorname{O}(V)$, $\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle=\sup_{q\in[z]}\langle x,q\rangle$ is a supremum of linear functionals.) In broad terms, an individual max filtering map $\langle\hskip-2.5pt\langle[\cdot],[z]\rangle\hskip-2.5pt\rangle$ is a scalar feature map which takes $x]\in V/G$ as input and measures the maximal alignment between the orbits $[x]$ and $[z]$ when interpreted as subsets of $V$. Notably, the map is a convex, $\|z\|$-Lipschitz continuous invariant, as it is defined as the supremum of $\|z\|$-Lipschitz linear functionals. A max filter bank consists of a collection of such maps, making it componentwise convex and Lipschitz continuous. Furthermore, it is uniquely determined up to origin-fixing isometries of $V/G$, as demonstrated by the following polarization identity, which holds for all $x,z\in V$:
$d([x],[z])^{2}=\|x\|^{2}-2\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle+\|z\|^{2},$ (1)
where we note that $\|x\|=d([x],[0])$.
The max filtering map can also be regarded as a fundamental convex invariant since every convex invariant $f\colon V\to\mathbb{R}$ can be expressed as a supremum of affine max filters, that is
$f(x)=\sup_{z\in\Omega}[\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle+b_{z}],$
for some $\Omega\subseteq V$ and $\{b_{z}\}_{z\in\Omega}\in\mathbb{R}^{\Omega}$. Additionally, when $G\leq\operatorname{O}(V)$ is compact, the subgradient of the max filtering map has an explicit form:
$\partial\langle\hskip-2.5pt\langle[\cdot],[z]\rangle\hskip-2.5pt\rangle|_{x}=\operatorname{conv}\{q\in[z]:\langle x,q\rangle=\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle\}.$
This follows from the general fact that $\partial(\max_{i\in I}f_{i})|_{x}=\operatorname{conv}\{\nabla f_{j}(x):j\in\arg\max_{i\in I}f_{j}(x)\}$, where $\{f_{i}\}_{i\in I}$ is a collection of convex differentiable functions and $\operatorname{conv}$ denotes the convex hull operator.
From a machine learning perspective, convexity is desirable because max filter banks can serve as $G$-invariant layers in classification models, with their constituent templates as trainable parameters. Relevant numerical examples are discussed in Section 6 of [9]. We also hypothesize that one could generalize input convex neural networks [4] to the convex invariant setting using max filters, but we leave this as a direction for future exploration.
A theoretical question posed in [9] asks whether every injective max filter bank is bilipschitz. When $G$ is finite, this question was resolved by [5], which showed that every injective max filter bank admits bilipschitz bounds. However, the question remains open for infinite $G$, with only three exceptions:
  • Complex phase retrieval [2, 7], where $V=\mathbb{C}^{d}$ and $G=\{z\cdot\operatorname{id}:z\in\mathbb{C},|z|=1\}$.
  • Polar actions [18], where $V/G$ is isometrically isomorphic to $V^{\prime}/G^{\prime}$ for some finite $G^{\prime}\leq\operatorname{O}(V^{\prime})$.
For general infinite $G$, the bilipschitz property of injective max filter banks remains an open question. While we do not fully resolve this issue, we investigate conditions under which these maps are bilipschitz, given sufficiently many generic templates. Specifically, we prove that this property holds locally near orbits of maximal dimension and hence globally for groups where all nonzero orbits have constant dimension.
1.1 Main results and paper outline
Our sufficient bound on the number of generic templates is determined by the following complexity parameter:
Definition 3.
Let $G\leq\operatorname{O}(d)$ be a compact group. The set of regular points is defined as
$R(G):=\big\{x\in\mathbb{R}^{d}:\dim([x])=\max_{y\in\mathbb{R}^{d}}\dim([y])\big\}.$
The regular Voronoi complexity of $G$ is then given by
$\chi(G):=\max_{x,p\in R(G)}\{|G_{x}/G_{p}|:G_{p}\leq G_{x}\},$
where $G_{y}:=\{g\in G:gy=y\}$ denotes the stabilizer of $y$ in $G$.
Note that the set $R(G)$ is an open, dense, $G$-invariant subset of $\mathbb{R}^{d}$ (for example, see Theorems 3.49 and 3.82 in [1]). Intuitively, $\chi(G)$ measures the maximum relative discrepancy in discrete degrees of freedom among orbits with the highest infinitesimal degrees of freedom, i.e., orbits of regular points.
We now present our two main results. The first establishes sufficient conditions under which max filter banks are locally bilipschitz over $R(G)$, without imposing restrictions on $G$ beyond compactness and orthogonality.
Theorem 4.
Let $G\leq\operatorname{O}(d)$ be a compact group and define $c:=d-\max_{x\in\mathbb{R}^{d}}\dim([x])$. For generic $z_{1},\dots,z_{n}\in\mathbb{R}^{d}$, the max filter bank $\Phi\colon\mathbb{R}^{d}/G\to\mathbb{R}^{n}$ given by $\Phi([x]):=\{\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle\}_{i=1}^{n}$ is locally bilipschitz at every $x\in R(G)$, provided $n>2\cdot\chi(G)\cdot(c-1)$.
Here, a map $f\colon\mathbb{R}^{d}/G\to\mathbb{R}^{n}$ is said to be locally bilipschitz at $x$ if it is bilipschitz when restricted to a neighborhood of $x$, with respect to the quotient distance induced by $G$. The above result is core to ensuring that max filter banks can form componentwise convex bilipschitz embeddings when all nonzero vectors lie within $R(G)$. This is the content of the following primary result:
Theorem 5.
Let $G\leq\operatorname{O}(d)$ be a compact group such that $\mathbb{R}^{d}-\{0\}\subseteq R(G)$, and define $c:=d-\max_{x\in\mathbb{R}^{d}}\dim([x])$. For generic $z_{1},\dots,z_{n}\in\mathbb{R}^{d}$, the max filter bank $\Phi\colon\mathbb{R}^{d}/G\to\mathbb{R}^{n}$ given by $\Phi([x]):=\{\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle\}_{i=1}^{n}$ is bilipschitz, provided $n>2\cdot\chi(G)\cdot(c-1)$.
In Section 1.2, we discuss the significance of Theorem 5 in the contexts of stable weighted phase retrieval and a nearest neighbor problem in cryo-EM. In Section 2, we introduce a Voronoi cell decomposition (Definition 8) that establishes a geometric framework for the proofs presented in this paper. We also give a key geometric characterization of $R(G)$ (Theorem 12) leveraging a desingularization theorem (Lemma 26) which may be of independent interest. The proofs of our main results (Theorems 4 and 5) are presented in Section 3. Finally, we conclude with a discussion in Section 4.
1.2 Relevance of results
In this section, we highlight the relevance of Theorem 5 in the contexts of stable weighted phase retrieval and a nearest neighbor problem in cryogenic electron microscopy (cryo-EM). Throughout, $\mathbb{C}^{*}\cong\operatorname{SO}(2)\cong S^{1}$ (resp. $\mathbb{H}^{*}\cong\operatorname{SU}(2)\cong S^{3}$) denotes the group of unit complex numbers (resp. unit quaternions).
1.2.1 Stable weighted phase retrieval
In this section, let $V$ be a finite dimensional real Hilbert space and let $G\leq\operatorname{O}(V)$ be a compact-connected group. The main result of this work (Theorem 5) establishes that max filter banks are bilipschitz if they include sufficiently many generic templates and if the nonzero orbits of $G$ have constant dimension.
In particular, if $\dim([p])=\dim(G)$ for some $p\in V$, the latter condition corresponds to the action of $G$ being almost free on the unit sphere $\mathbb{S}(V)$, meaning that every nonzero orbit has a finite stabilizer. As noted in Section 3.2 in [13] and since $G$ is connected, this happens if and only if $G$ is the image of a representation $\phi\colon K^{*}\to\operatorname{O}(V)$, where $K=\mathbb{C}$ or $K=\mathbb{H}$, and $\phi$ is a direct sum of nontrivial irreducible complex (resp. quaternionic) representations of $K^{*}$. In this case, we call $\mathbb{S}(V)/G$ a weighted complex (resp. quaternionic) projective space.
In these settings, templates $z_{1},\dots,z_{n}\in V$ are said to achieve stable weighted phase retrieval if the max filter bank $\Phi\colon V/G\to\mathbb{R}^{n}$, defined by $\Phi([x])=\{\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle\}_{i=1}^{n}$, is bilipschitz. This leads to the following corollary of Theorem 5:
Corollary 6.
Suppose that $G\leq\operatorname{O}(V)$ is compact, connected and acting almost freely on $\mathbb{S}(V)$. Then, $\chi(G)=\max_{x\neq 0}|G_{x}|$ and generic templates $z_{1},\dots,z_{n}\in V$ achieve stable weighted phase retrieval provided $n>2\cdot\chi(G)\cdot(c-1)$, where $c:=\dim_{\mathbb{R}}(V)-\dim(G)$.
This framework generalizes stable complex phase retrieval [2, 7], where one seeks to stably recover $x\in\mathbb{C}^{d}$, up to the equivalence $x\sim e^{\mathrm{i}\theta}x$, from magnitude measurements $|x^{*}z_{i}|=\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle$. This setup corresponds to a bilipschitz max filter bank.
In the complex setting, stable weighted phase retrieval generalizes this by incorporating weights into the orbit equivalence relation induced by $G\leq\operatorname{O}(\mathbb{C}^{d})$:
$x\sim\operatorname{diag}\{e^{\mathrm{i}k_{j}\theta}\}_{j=1}^{d}\cdot x,\quad\forall\,\theta\in[-\pi,\pi],$
where $\{k_{j}\}_{j=1}^{d}$ is a collection of nonzero integers, namely the weights. Viewing this problem from the lens of $\phi\colon K^{*}\to O(V)$ as above, we obtain the natural extension to the quaternionic setting.
1.2.2 Approximate nearest neighbor problem in cryo-EM
Given a 3D macromolecular complex $P$ with unknown structure, cryo-EM produces a set of 2D noisy projection images $\mathcal{I}=\{I_{j}\}_{j=1}^{M}$, corresponding to different 3D viewing angles. As argued in [19, 20], each image $I_{j}$ can be approximated as an $L\times L$ pixel sampling of a function $f_{j}\in L^{1}(\mathbb{R}^{2})\cap L^{2}(\mathbb{R}^{2})$, with Fourier transform $\mathcal{F}f_{j}\in L^{2}(\mathbb{R}^{2})$ compactly supported in the disk $\Omega_{c_{0}}\subseteq\mathbb{R}^{2}$ of radius $c_{0}$. Furthermore, $\mathcal{F}f_{j}$ is approximated in polar coordinates by a finite expansion
$\mathcal{F}f_{j}(r,\theta)\approx\sum_{k=-k_{\max}}^{k_{\max}}\sum_{q=1}^{p_{k}}a_{k,q}^{f_{j}}\psi_{k,q}^{c_{0}}(r)e^{ik\theta},$ (2)
where $\{\psi_{k,q}^{c_{0}}(r)e^{ik\theta}\}_{k\in\mathbb{Z},q\in\mathbb{N}}$ is the scaled Fourier-Bessel $L^{2}(\Omega_{c_{0}})$-orthonormal basis, and one may take $k_{max}=O(L)$ and $p:=\sum_{-k_{\max}}^{k_{\max}}p_{k}=O(L^{2})$.
Denoising an image $I_{j}$ involves finding its nearest neighbors under the rotational alignment distance:
$d_{\operatorname{SO}(2)}(f,g):=\inf_{R\in\operatorname{SO}(2)}\big\|f-g\circ R^{-1}\big\|_{2}.$
Since this is computationally expensive, an approximation via (2) is used:
$d_{\operatorname{SO}(2)}(I_{i},I_{j})\approx\min_{\alpha\in[0,2\pi)}\big\|\big\{a_{k,q}^{f_{i}}\big\}-\big\{a_{k,q}^{f_{j}}e^{\mathrm{i}k\alpha}\big\}\big\|_{2}.$
This shifts the setting of the nearest neighbor task into the orbit space of a finite dimensional orthogonal representation $W$ of $\mathbb{C}^{*}$ given by the space of coefficients
$W:=\big\{\{a_{k,q}\}:|k|\leq k_{\max},p_{k}>0,q\in\{1,\dots,p_{k}\}\big\},$
equipped with the unitary action $z\cdot\{a_{k,q}\}:=\{z^{k}\cdot a_{k,q}\}$, for $z\in\mathbb{C}^{*}$. By virtue of Example 1, we seek a fast bilipschitz embedding of $W/\mathbb{C}^{*}$ into Euclidean space. In [19], the authors use the bispectrum embedding $\Phi_{B}\colon W/\mathbb{C}^{*}\to\mathbb{R}^{N}$, with $N=O(L^{5})$, defined by
$\Phi_{B}(\mathbb{C}^{*}\cdot\{a_{k,q}\}):=\{a_{k_{1},q_{1}}a_{k_{2},q_{2}}\overline{a_{k_{1}+k_{2},q_{3}}}:p_{k_{1}},p_{k_{2}},p_{k_{1}+k_{2}}>0\}.$
However, by Proposition 41 in Appendix C, $\Phi_{B}$ fails to be bilipschitz if $p_{k},p_{k^{\prime}}>0$ for some $k\neq\pm k^{\prime}$. The authors in [19] also suggest precomposing $\Phi_{B}$ with the continuous scaling $\{a_{k,q}\}\mapsto\left\{{a_{k,q}}/{|a_{k,q}|^{2/3}}\right\}$, but this fails to be upper Lipschitz.
We propose a replacement by using max filter banks. By projecting away from the trivial component as in Remark 7 below, the problem reduces to stable weighted phase retrieval. Then Corollary 6 applies with $\chi(G)\leq k_{\max}=O(L)$ and $c\leq p=O(L^{2})$, and it entails that $O(L^{3})$ generic templates constitute bilipschitz max filter banks (cf. $N=O(L^{5})$ above.) Moreover, each constituent max filtering map can be approximated by a linear search over samples obtained via the fast Fourier transform. We hypothesize that max filter banks will outperform the bispectrum in accuracy and efficiency, but since our focus is theoretical, we leave the numerical validation of this for future work.
Remark 7.
Suppose $G\leq\operatorname{O}(W)$ is compact and let $F:=\{w\in W:G_{w}=G\}$ denote its fixed subspace. Put $V:=F^{\perp}$ and let $P_{F}$ and $P_{V}$ denote corresponding linear orthogonal projections. By the proof of Lemma 39 in [8] and given a bilipschitz embedding $f\colon V/G\to\mathbb{R}^{n}$ with bounds $\alpha\leq\beta$, the map $\Psi\colon W/G\to F\times\mathbb{R}^{n}$ defined by $\Psi([x]):=\big(\alpha\cdot P_{F}x,\beta\cdot f([P_{V}x])\big)$ also has bilipschitz bounds $\alpha\leq\beta$.
2 The Voronoi Decomposition
In this section, we conduct a geometric analysis of isometric linear actions by introducing a Voronoi cell decomposition of the space. A key result, Theorem 12, will play a central role in proving the key Lemma 30 of Section 3.
We begin in Section 2.1 by introducing the decomposition and presenting the main results of this section: Theorems 10 and 12. In Section 2.2, we provide a concrete example to illustrate the underlying concepts. Section 2.3 covers the necessary preliminaries, while Section 2.4 establishes and proves key properties of the Voronoi decomposition. Finally, Sections 2.5 and 2.6 contain the proofs of Theorems 10 and 12, respectively.
2.1 Setup and main results
For a subset $S\subseteq\mathbb{R}^{d}$, let $\operatorname{relint}(S)$ denote the interior of $S$ relative to its affine span $\operatorname{aff}(S)$.
Definition 8.
Let $G\leq\operatorname{O}(d)$ be a compact group. For $x\in\mathbb{R}^{d}$, the unique Voronoi cell of $x$, denoted $U_{x}$, is defined by:
$z\in U_{x}\iff\{x\}=\arg\max_{p\in[x]}\langle p,z\rangle\iff\{x\}=\arg\min_{p\in[x]}\|z-p\|.$
The open Voronoi cell of $x$ is then defined as:
$V_{x}:=\operatorname{relint}\big(U_{x}\big),$
and the open Voronoi diagram of $x$ is given by
$Q_{x}:=\bigsqcup_{p\in[x]}V_{p}.$
The equivalence of the two characterizations of $U_{x}$ follows directly from the polarization identity $\|z-p\|^{2}=\|x\|^{2}+\|z\|^{2}-2\langle p,z\rangle$ for $p\in[x]$. In Section 2.4, we explore the properties of this decomposition and provide characterizations for its components.
The distinction between $U_{x}$ and its relative interior $V_{x}$ is important. First, the definition is non-redundant, as $U_{x}\neq V_{x}$ may occur (e.g., see Example 20.) Second, $V_{x}$ is functionally essential, as demonstrated in the proof of Lemma 30 in Section 3. There, we use Theorem 12, which provides a geometric “interchangeability” characterization of $V_{x}$ when $[x]$ has maximal dimension, i.e., $x\in R(G)$. To state both of our Voronoi ‘interchangeability’ results, we first introduce the concept of principality.
Definition 9.
Let $G\leq\operatorname{O}(d)$ be compact. The set of principal points is defined as
$P(G):=\big\{x\in\mathbb{R}^{d}:\forall\,z\in\mathbb{R}^{d},G_{z}\leq G_{x}\implies G_{z}=G_{x}\}.$
Notably, $P(G)\subseteq R(G)$ is a $G$-invariant open and dense subset of $\mathbb{R}^{d}$, as established in Theorems 3.49 and 3.82 in [1]. Intuitively, orbits of principal points possess maximal degrees of freedom, both in the infinitesemal and discrete sense.
In Section 2.5, we prove the following geometric characterization of principal points in terms of open Voronoi cells.
Theorem 10.
Let $G\leq\operatorname{O}(d)$ be compact. The following are equivalent:
  • (a)
    $x\in P(G)$.
  • (b)
    $z\in V_{x}$ implies $x\in V_{z}$.
Proof.
See Section 2.5. ∎
Building on this, we refine the Voronoi decomposition to the local level, enabling a similar characterization for regular points.
Definition 11.
Let $G\leq\operatorname{O}(d)$ be compact. For $z\in\mathbb{R}^{d}$, the local open Voronoi cell $V_{z}^{loc}$, is defined as follows: $x\in V_{z}^{loc}$ if there exist open neighborhoods $U$ of $z$ and $V$ of $x$ such that:
$\forall q\in V,\ \ \Big|\arg\sup_{p\in[z]\cap U}\langle p,q\rangle\Big|=1.$
Theorem 12.
Let $G\leq\operatorname{O}(d)$ be compact. For $x\in\mathbb{R}^{d}$, the following are equivalent:
  1. (a)
    $x\in R(G)$.
  2. (b)
    $z\in V_{x}$ implies $x\in V_{z}^{loc}$.
Proof.
See Section 2.6. ∎
Although Theorem 10 is not directly referenced later in the paper, it is an independently interesting geometric result. Additionally, its proof in Section 2.5 provides preparatory groundwork for the more technical and analogous proof of Theorem 12 in Section 2.6.
2.2 Concrete example
In this section, we provide a concrete three-dimensional example to illustrate the definitions and theorems we have presented in this section so far. In this example, we will observe that the action behaves ‘nicely’ in the sense that $U_{x}=V_{x}$ for all $x\in\mathbb{R}^{3}$. However, as we will see later in Example 20, this is not always the case.
Figure 1: Illustration for Example 13, showing instances of $x$, $y$, and $z$ along with their orbits and Voronoi cells, all of which are precisely described in the referenced example.
Example 13.
Suppose $G\leq\operatorname{O}(3)$ is the commutative group generated by $\operatorname{SO}_{Z}(2)$, the subgroup consisting of all counterclockwise rotations around the $Z$-axis, and $R_{XY}$, the reflection across the $XY$-plane. We begin by computing $U_{w}$ and $V_{w}^{loc}$ for each $w\in\mathbb{R}^{3}$, dividing analysis into cases; see Figure 1.
For $x\in\{(0,0,x_{3})\in\mathbb{R}^{3}:x_{3}>0\}$, it holds that $V_{x}^{loc}=\mathbb{R}^{3}$ since all points project uniquely onto $\{x\}$. Next, $U_{x}$ is the open upper half-space (depicted in blue in Figure 1.) Similarly, $U_{R_{XY}x}=R_{XY}U_{x}$. This covers the case of nonregular nonzero points.
For $y\in\{(0,x_{2},x_{3})\in\mathbb{R}^{3}:x_{2},x_{3}>0\}$, it holds that $V_{y}^{loc}$ is the complement of the $Z$-axis, as one can take a tubular neighborhood $U$ of $\operatorname{SO}_{Z}(2)\cdot y$ so that all points, except those lying on the $Z$-axis, project uniquely onto $y]\cap U=\operatorname{SO}_{Z}(2)\cdot y$. Next, $U_{y}$ is the open first quadrant of the $YZ$-plane (shown in orange in Figure 1.) Additionally, $U_{gy}=gU_{y}$ for all $g\in G$. This covers the case of principal points.
For $z\in\{(0,-x_{2},0)\in\mathbb{R}^{3}:x_{2}>0\}$, we have that $V_{z}^{loc}$ is again the complement of the $Z$-axis by a similar argument as in the previous paragraph. Next, $U_{z}$ is given by the open half-plane $\{(0,-x_{2},x_{3})\in\mathbb{R}^{3}:x_{2}>0,x_{3}\in\mathbb{R}\}$. Additionally, $U_{R_{XY}z}=U_{z}$ and $U_{gz}=gU_{z}$ for each $g\in G$. This covers the case of regular nonprincipal points.
Lastly, $U_{(0,0,0)}=\mathbb{R}^{3}$ and $V_{(0,0,0)}^{loc}=\mathbb{R}^{3}$.
Note that $V_{w}=U_{w}$ for each $w\in\mathbb{R}^{3}$ since in each of the above cases, $U_{w}$ is open in its affine hull. Moreover, $Q_{x}$ is the complement of the $XY$-plane, $Q_{z}$ is the complement of the $Z$-axis, $Q_{y}=Q_{x}\cap Q_{z}$, and $Q_{(0,0,0)}=\mathbb{R}^{3}$.
[…]
2.3 Preliminary Results
We begin with a crucial preliminary result drawn from the theory of nonlinear orthogonal projection on manifolds, which will be referenced frequently throughout the paper. A visual illustration is provided in Figure 2. For $x\neq y\in\mathbb{R}^{d}$, let $(x,y]$ denote the line segment from $x$ to $y$, which includes $y$ but excludes $x$, and let $[x,y]$ denote the segment that includes both $x$ and $y$; additionally, define $(x,x]:=[x,x]=\{x\}$.
Proposition 14 (Remark 3.1, Corollary 3.9, Theorem 3.13a and Theorem 4.1 in [11]).
Let $M$ be a smooth embedded submanifold of $\mathbb{R}^{d}$. For $z\in\mathbb{R}^{d}$ and $x\in\arg\min_{p\in M}\|z-p\|$, each of the following statements holds:
  1. (a)
    $z\in N_{x}M$, the orthogonal complement of the tangent space to $M$ at $x$.
  2. (b)
    There exists an open neighborhood $U$ of $(z,x]$ such that $\{x\}=\arg\min_{p\in M}\|n-p\|$ for $n\in U\cap N_{x}M$ and $\big|\arg\min_{p\in M}\|t-p\|\big|=1$ for all $t\in U$. Moreover, the map $v_{x}$, which sends $t\in U$ to the unique element in $\arg\min_{p\in M}\|t-p\|$, is smooth over $U$.
  3. (c)
    If there exists an open neighborhood $W_{z}$ around $z$ such that $|\arg\min_{p\in M}\|t-p\||=1$ for all $t\in W_{z}$, then the neighborhood $U$ in (b) can be enlarged to include $z$.
Remark 15.
If $M\subseteq\mathbb{R}^{d}$ lies on a sphere centered at the origin (e.g., $M=G\cdot x$ for some $G\leq\operatorname{O}(d)$), then for all $z\in\mathbb{R}^{d}$, the polarization identity gives
$\arg\min_{p\in M}\|z-p\|=\arg\max_{p\in M}\langle p,z\rangle.$
Figure 2: Illustration for Proposition 14. (left) Here, $M=\{(t,t^{2})\in\mathbb{R}^{2}:t\in\mathbb{R}\}$ and $U=\{(x_{1},x_{2})\in\mathbb{R}^{2}:x_{1}>0\}$. The line $L=\{(0,x_{2})\in\mathbb{R}^{2}:x_{2}>\frac{1}{2}\}$ is the set of points with multiple nearest neighbors to $M$, and its closure is $\overline{L}=L\cup\{(0,\frac{1}{2})\}$. For each $z\in U\cup L$ and $x\in U\cap\arg\min_{p\in M}\|p-z\|$, the neighborhood $U$ satisfies assertions (b) and (c) in Proposition 14. For $z_{0}\in R:=\{(0,x_{2})\in\mathbb{R}^{2}:x_{2}<\frac{1}{2}\}$, it holds that $\arg\min_{p\in M}\|p-z\|=\{(0,0)\}$, and any neighborhood $U$ of $R$ with $U\cap\overline{L}=\varnothing$ satisfies those assertions with respect to each $z_{0}\in R$. (right) Here, $M=\{(\cos(\theta),\sin(\theta))\in\mathbb{R}^{2}:\theta\in(-\frac{\pi}{3},\frac{\pi}{3})\}$ and $U=\operatorname{int}(\operatorname{cone}(M))$. For each $z\in U\cup\{(0,0)\}$ and $x\in\arg\min_{p\in M}\|p-z\|$, the neighborhood $U$ satisfies assertions (b) and (c) in Proposition 14. For nonzero $z\in U^{c}$, the set $\arg\min_{p\in M}\|p-z\|$ is empty.
Next, for a compact group $G\leq\operatorname{O}(d)$ with Lie algebra $\mathfrak{g}$ and $x\in\mathbb{R}^{d}$, the orbit $G\cdot x$ is an embedded submanifold of $\mathbb{R}^{d}$ (Proposition 3.41 in [1]). We define the tangent space at $x$ to its orbit $G\cdot x$ by
$T_{x}:=\mathfrak{g}\cdot x,$
and the normal space at $x$ to its orbit $G\cdot x$ as:
$N_{x}:=(\mathfrak{g}\cdot x)^{\perp}.$
We have the following (presumably folklore) result which establishes the $G_{x}$-invariance of $T_{x}$ and $N_{x}$. While we could not locate a reference, we provide a proof.
Proposition 16.
Let $G\leq\operatorname{O}(d)$ be a compact group and fix $x\in\mathbb{R}^{d}$. Then, for each $h\in G$, we have $T_{hx}=h\cdot T_{x}$ and $N_{hx}=h\cdot N_{x}$. In particular, $T_{x}\oplus N_{x}$ is a $G_{x}$-invariant orthogonal decomposition of $\mathbb{R}^{d}$.
Proof.
Since each $h\in G$ is an isometry, it suffices to show that $T_{hx}=h\cdot T_{x}$; this would then imply that $h\cdot N_{x}=h\cdot(T_{x})^{\perp}=(T_{hx})^{\perp}=N_{hx}$. Since $h$ is linear and $\dim(T_{x})=\dim(T_{hx})$, it suffices to show that $h\cdot T_{x}\subseteq T_{hx}$ for each $h\in G$. To this end, fix $h\in G$ and $t\in T_{x}$. We aim to show that $ht\in T_{hx}$. By the definition of $T_{x}$, there exists $\omega\in\mathfrak{g}$ such that $t=\omega\cdot x$. Since $\mathfrak{g}$ is the tangent space to $G$ at its identity $e$, there exists $\varepsilon>0$ and a smooth curve $\alpha\colon(-\varepsilon,\varepsilon)\to G$ such that $\alpha(0)=e$ and $\alpha^{\prime}(0)=\omega$. The smoothness of the action of $G$ on $\mathbb{R}^{d}$ implies:
$t=\omega\cdot x=\alpha^{\prime}(0)\cdot x=\left.\frac{d(\alpha(t)x)}{dt}\right|_{t=0}.$
Now, let $\gamma:=h\alpha h^{-1}\colon(-\varepsilon,\varepsilon)\to G$ denote the conjugation of $\alpha$ by $h$. Then $\gamma$ is a smooth curve satisfying $\gamma(0)=e$ and $\gamma^{\prime}(0)\in\mathfrak{g}$. Thus
$ht=h\cdot\left.\frac{d(\alpha(t)x)}{dt}\right|_{t=0}=\left.\frac{d(h\alpha(t)h^{-1}(hx))}{dt}\right|_{t=0}=\gamma^{\prime}(0)\cdot hx\in T_{hx}.$
2.4 Properties of the Voronoi Decomposition
The following lemma explores the properties of $U_{x}$ and $V_{x}$, and the third statement provides justification for using a disjoint union in the definition of $Q_{x}$.
Lemma 17.
Let $G\leq\operatorname{O}(d)$ be a compact group. For $x\in\mathbb{R}^{d}$, each of the following statements holds:
  1. (a)
    $z\in U_{x}$ if and only if $G_{x}=\{g\in G:\langle\hskip-2.5pt\langle[z],[x]\rangle\hskip-2.5pt\rangle=\langle gz,x\rangle\}$.
  2. (b)
    $U_{gx}=g\cdot U_{x}$ and $V_{gx}=g\cdot V_{x}$ for all $g\in G$.
  3. (c)
    For $q_{1},q_{2}\in[x]$, if $U_{q_{1}}\cap\overline{U_{q_{2}}}\neq\varnothing$, then $q_{1}=q_{2}$.
  4. (d)
    $\operatorname{aff}(U_{x})=\operatorname{span}(U_{x})=N_{x}$, and $V_{x}$ is a star convex open neighborhood of $x$ in $N_{x}$.
  5. (e)
    The following characterization holds:
[…]
  6. (f)
    The sets $N_{x}$, $U_{x}$ and $V_{x}$ are semialgebraic subsets of $\mathbb{R}^{d}$.
Proof.
The proofs of (a), (b) and (c) are straightforward.
To prove (d), note that $U_{x}\subseteq N_{x}$ by 14(a), so we have $\operatorname{aff}(U_{x})\subseteq\operatorname{span}(U_{x})\subseteq N_{x}$. Since $x\in U_{x}$, 14(b) guarantees the existence of an open neighborhood $U$ of $(x,x]=\{x\}$ in $\mathbb{R}^{d}$ such that $U\cap N_{x}$ is a nonempty subset of $U_{x}$. This implies that $N_{x}=\operatorname{aff}(U\cap N_{x})\subseteq\operatorname{aff}(U_{x})$, so $\operatorname{aff}(U_{x})=\operatorname{span}(U_{x})=N_{x}$. Furthermore, $U$ is witnesses that $V_{x}$ is open in $N_{x}$ and that $x\in V_{x}$. For the star convexity of $V_{x}$ at $x$, let $z\in V_{x}$ and note that by 14(b), there exists an open neighborhood $U^{\prime}$ of $(z,x]$ such that $(z,x]\subseteq U^{\prime}\cap N_{x}\subseteq V_{x}$.
For the forward implication in (e), the openness of $V_{x}$ in $N_{x}$ and its star convexity at $x$ imply that there exists $q\in V_{x}$ such that $z,x]\subseteq(q,x]\subseteq V_{x}$. Then the desired result follows from 14(b) applied to the interval $(q,x]$. Next, the reverse implication in (e) follows immediately from 14(b) and 14(c) applied to the interval $[z,x]$.
Lastly, (f) follows from a straightforward argument in first-order logic. A restatement and proof can be found in Proposition 40. ∎
As a first application of Lemma 17, we derive stabilizer inclusions implied by the Voronoi decomposition.
Lemma 18.
Let $G\leq\operatorname{O}(d)$ be a compact group. For $x,z\in\mathbb{R}^{d}$, each of the following statements holds:
  1. (a)
    $z\in U_{x}$ implies $G_{z}\leq G_{x}$.
  2. (b)
    $x\in P(G)$ and $z\in N_{x}$ imply $G_{x}\leq G_{z}$.
  3. (c)
    $x\in P(G)$ and $z\in U_{x}$ imply $G_{x}=G_{z}$ and $z\in P(G)$.
Intuitively, (a) states that if $z$ uniquely projects to $x$ within $[x]$, then $z$ has $G_{x}/G_{z}$’ more degrees of freedom than $x$. (b) states that if $x\in P(G)$, meaning $x$ has maximal degrees of freedom among all orbits, then $N_{x}$ is fixed by $G_{x}$, i.e., while fixing $x$, $G_{x}$ does not introduce any degrees of freedom to $N_{x}$. Finally, (c) combines (a) and (b).
Proof of Lemma 18.
First, we address (a). For $z\in U_{x}$ and $g\in G_{z}$, we have $z=gz\in U_{x}$. By 17(b), we have $z\in U_{g^{-1}x}\cap U_{x}$, and by 17(c), it follows that $g\in G_{x}$, as required.
Next, we prove (b). For each $y\in U_{x}$, (a) gives that $G_{y}\leq G_{x}$. Since $x\in P(G)$, it follows that $G_{y}=G_{x}$. Thus, $G_{x}$ fixes $U_{x}$, and by linearity and 17(d), we conclude that $G_{x}$ fixes $\operatorname{span}(U_{x})=N_{x}$.
Finally, (c) follows immediately from (a) and (b). ∎
The following lemma elaborates on the properties of the open Voronoi diagram $Q_{x}$.
Lemma 19.
Let $G\leq\operatorname{O}(d)$ be a compact group. For $x\in\mathbb{R}^{d}$, each of the following statements holds:
  1. (a)
    $Q_{gx}=Q_{x}=g\cdot Q_{x}=G\cdot V_{x}$ for all $g\in G$.
  2. (b)
    The set $Q_{x}$ is a semialgebraic subset of $\mathbb{R}^{d}$.
  3. (c)
    The following characterization holds:
[…]
  4. (d)
    $Q_{x}$ is an open and dense subset of $\mathbb{R}^{d}$.
Proof.
The proofs of (a) is straightforward.
Part (b) follows from a standard argument using first-order logic. A restatement and proof are provided in Proposition 40.
The proof of (c) directly follows from 17(e).
Finally, we address (d). The openness of $Q_{x}$ follows immediately from the openness of the characterization in (c). For denseness, let $z\in\mathbb{R}^{d}$ and choose $q\in\arg\max_{p\in[x]}\langle p,z\rangle$, which is possible because $[x]$ is closed. By 14(b), there exists an open neighborhood $U$ of $(z,q]$ such that $|\arg\max_{p\in[x]}\langle p,u\rangle|=1$ for all $u\in U$. By (c), we have $U\subseteq Q_{x}$. Since $z\in\overline{(z,q]}\subseteq\overline{U}\subseteq\overline{Q_{x}}$, the denseness of $Q_{x}$ follows. ∎
As a first application of Lemma 19, we identify the unitary representations of the circle group for which $U_{x}=V_{x}$ for all $x\in\mathbb{R}^{d}$.
Example 20.
Let $k_{1},\dots,k_{d}\in\mathbb{Z}$ be fixed integers, and let $G\leq\operatorname{U}(d)$ be the commutative group defined by $G:=\{\operatorname{diag}(\{e^{\mathrm{i}k_{j}\theta}\}_{j=1}^{d}):\theta\in[-\pi,\pi]\}$. We compute the max filter and then perform a case analysis on $\{k_{j}\}_{j=1}^{d}$ to determine when $U_{x}=V_{x}$ for all $x\in\mathbb{R}^{d}$.
Suppose, without loss of generality, that $\{k_{j}\}_{j=1}^{d}$ is sorted in ascending order. Then there exist $l\in\mathbb{N}$, $d_{1},\dots,d_{l}\in\mathbb{N}$, and distinct weights $w_{1},\dots,w_{l}\in\mathbb{Z}$ such that $k_{j}=w_{m}$ for $j\in r_{m}:=\{d_{m-1}+1,\dots,d_{m}\}$ (with $d_{0}:=0$.) For $x\in\mathbb{R}^{d}$, we have an orthogonal decomposition $x=\sum_{m=1}^{l}x_{m}$, where $(x_{m})_{j}:=1_{j\in r_{m}}\cdot(x)_{j}$ for each $j\in\{1,\dots,d\}$.
With this notation, the max filter is given by
$\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle=\max_{-\pi\leq\theta\leq\pi}\,\operatorname{Re}\left(\sum_{m=1}^{l}x_{m}^{*}z_{m}e^{\mathrm{i}w_{m}\theta}\right).$
We claim that $U_{x}=V_{x}$ holds for all $x\in\mathbb{R}^{d}$ if and only if $l=1$ or $l=2$ with $0\in\{w_{1},w_{2},w_{1}+w_{2}\}$. These cases correspond to the nontrivial component of the action having a spherical quotient diffeomorphic to a standard complex projective space.
Case 1. Suppose that $l=2$ with $0\in\{w_{1},w_{2}\}$. This case reduces immediately to $l=1$. For if $w_{1}=0$, then it holds that $U_{x}=\mathbb{C}^{d_{1}}\times U_{x_{2}}$ and $V_{x}=\mathbb{C}^{d_{1}}\times V_{x_{2}}$ where $U_{x_{2}}$ and $V_{x_{2}}$ correspond to the case $l=1$ and unique weight $w_{2}$.
Case 2. Suppose that $l=2$ with $w_{1}=-w_{2}$. This case is orthogonally equivalent the case $l=1$ with unique weight $w_{1}$. This can be seen by the map $x=x_{1}+x_{2}\mapsto x_{1}+\overline{x_{2}}$, which is an orthogonal (nonunitary) transformation of space.
Case 3. Suppose that $l=1$ and $x\neq 0$ (indeed, $U_{0}=V_{0}=\mathbb{C}^{d}$.) Then
$\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle=\max_{-\pi\leq\theta\leq\pi}\,\operatorname{Re}\left(x^{*}ze^{\mathrm{i}w_{1}\theta}\right)=|x^{*}z|.$
If $x^{*}z=0$, then $(gx)^{*}z=0$ for each $g\in G$ implying $z\notin G\cdot U_{x}$. Otherwise, if $x^{*}z=re^{\mathrm{i}\phi}$, the condition $\operatorname{Re}(re^{\mathrm{i}\phi}e^{\mathrm{i}w_{1}\theta})=r$ holds if and only if $\phi+w_{1}\theta\in 2\pi\mathbb{Z}$, which uniquely determines $e^{\mathrm{i}w_{1}\theta}$. This implies that $z\in G\cdot U_{x}$. As such
$G\cdot U_{x}=\{z\in\mathbb{R}^{d}:x^{*}z\neq 0\}.$
Since $x^{*}z=0$ is a closed condition, it follows that $G\cdot U_{x}$ is open and hence equal to $Q_{x}=G\cdot V_{x}$ by 19(c). By 17(c), we conclude that $U_{x}=U_{x}\cap G\cdot U_{x}=U_{x}\cap G\cdot V_{x}=V_{x}$ for all $x\in\mathbb{R}^{d}$.
Figure 3: Illustration for Case 4 in Example 20 when $w_{1}=1$ and $w_{2}=2$. The graph with the highest $y$-intercept of 33 corresponds to the function $y=4\cos(x)-\cos(2x)$. It has a unique global maximum over $\pi,\pi]$, attained as a flat local maximum at $x=0$. The other graphs, with $y$-intercepts k1k-1, correspond to the functions $y=k\cos(x)-\cos(2x)$ for $k\in\{1,2,3,3.5\}$. Each attains its global maximum over $\pi,\pi]$ at $x=\pm\cos^{-1}\left(k/4\right)\neq 0$. While the values of $k$ here do not exactly correspond to $4-\frac{1}{n}$, the behavior of the global maxima remains the same for that sequence. (We thank Aleksei Kulikov for bringing this example to our attention.)
Case 4. Suppose that $l\geq 2$ and $w_{1},w_{2},w_{1}+w_{2}\neq 0$. Define $x\in\mathbb{R}^{d}$ and a sequence $z_{n}\in\mathbb{R}^{d}$ as follows:
$\begin{array}[]{cc}(x)_{j}=\begin{cases}1&\text{if }j\in\{1,d_{1}+1\},\\ 0&\text{otherwise,}\end{cases}&\text{\quad and \quad}(z_{n})_{j}=\begin{cases}w_{2}^{2}-\frac{1}{n}&\text{if }j=1,\\ -w_{1}^{2}&\text{if }j=d_{1}+1,\\ 0&\text{otherwise.}\end{cases}\end{array}$
Then
$\langle\hskip-2.5pt\langle[x],[z_{n}]\rangle\hskip-2.5pt\rangle=\max_{-\pi\leq\theta\leq\pi}\left(w_{2}^{2}-\frac{1}{n}\right)\cos(w_{1}\theta)-w_{1}^{2}\cos(w_{2}\theta).$
Let $z:=\lim_{n\to\infty}z_{n}$. Then $z\in U_{x}$, but $z_{n}\notin Q_{x}$ for each $n$ (see Figure 3.) Since $Q_{x}$ is open (19(d)), we have $z\in U_{x}\cap Q_{x}^{c}=U_{x}\cap V_{x}^{c}$, which means that $U_{x}\neq V_{x}$, as desired.
2.5 Geometric Characterization of Principal Orbits
This section is dedicated to proving Theorem 10. To enhance readability and help absorb the core ideas, we recommend skipping the proofs of Propositions 22, 25 and 23 on the first read. To appreciate the nuance in Theorem 10, consider the following easier characterization of principality.
Lemma 21.
Suppose $G\leq\operatorname{O}(d)$ is compact. For $x\in\mathbb{R}^{d}$, the following statements are equivalent:
  1. (a)
    $x\in P(G)$.
  2. (b)
    $z\in U_{x}$ implies $x\in U_{z}$.
Proof.
[…]
$G_{z}=G_{x}^{-1}=\{g^{-1}\in G:\langle\hskip-2.5pt\langle[z],[x]\rangle\hskip-2.5pt\rangle=\langle gz,x\rangle\}=\{g\in G:\langle\hskip-2.5pt\langle[x],[z]\rangle\hskip-2.5pt\rangle=\langle gx,z\rangle\}.$
Thus, another application of 17(a) shows that $x\in U_{z}$.
[…]
Proposition 22.
Suppose $G\leq\operatorname{O}(d)$ is compact. Then, each of the following statements holds
  • (a)
    $[P(G)]$ is an open and dense connected manifold in $\mathbb{R}^{d}/G$. It admits a unique Riemannian structure, whose geodesic distance agrees with the quotient distance, and where the restricted orbit map $\cdot]\big|_{P(G)}\colon P(G)\to[P(G)]$ is a Riemannian submersion.
  • (b)
    For each $x,y\in P(G)$, let $C([x],[y])$ denote the set of minimal geodesics joining $[x]$ to $[y]$ in $[P(G)]$. Then, there exists a bijection
    $\arg\min_{q\in[y]}\|q-x\|\longrightarrow C([x],[y])$
    induced by projecting straight lines joining $x$ to $\arg\min_{q\in[y]}\|q-x\|$ into the orbit space.
Proof.
For (a), the openness, denseness and connectedness of $[P(G)]$ follow from Theorem 3.82 in [1]. The rest of the statement regarding the unique Riemannian structure follows from Exercise 3.81 in [1].
For (b), Kleiner’s Lemma (Lemma 3.70 in [1]) and 14(a) imply that straight lines joining $x$ to $\arg\min_{q\in[y]}\|q-x\|$ lie entirely in $P(G)$ and $N_{x}$, respectively. In particular, 18(b) implies that every element of $\arg\min_{q\in[y]}\|q-x\|$ is fixed by $G_{x}$. Then, Lemma 40 in [18] yields the desired bijection
$\arg\min_{q\in[y]}\|q-x\|=\frac{\arg\min_{q\in[y]}\|q-x\|}{G_{x}}\longrightarrow C([x],[y]),$
induced by projecting (unit speed) straight lines joining $x$ to $\arg\min_{q\in[y]}\|q-x\|$ into the orbit space wherein they land in $[P(G)]$. ∎
Next, we reformulate the implication (a) $\Rightarrow$ (b) in Theorem 10 using a ‘cut-locus’ interchangeability argument for the Riemannian manifold $[P(G)]$, via the following lemma.
Lemma 23.
Suppose $G\leq\operatorname{O}(d)$ is compact. For $x,z\in P(G)$, the following statements are equivalent:
  1. (a)
    $z\in Q_{x}$.
  2. (b)
    There exists $q\in P(G)$ and a minimal geodesic $\gamma^{r}$ in $[P(G)]$ joining $[x]$ to $[q]$ such that $[z]$ lies in the interior of $\gamma^{r}$.
Proof.
[…]
If the geodesic metric of $[P(G)]$ were complete, the last statement of Lemma 23 would be symmetric in $[x]$ and $[z]$. This is due to the remarkable symmetry of the cut locus for metrically complete Riemannian manifolds (Scholium 3.78 in [12].) Although $[P(G)]$ is almost never metrically complete, we are still able to prove the desired symmetry for all Riemannian manifolds with the help of additional hypotheses which $[P(G)]$ satisfies. Since the proof is an adaptation of the arguments for Proposition 13.2.2 in [10], we postpone it to the appendix.
Lemma 24.
Let $M$ be a Riemannian manifold and let $x],[z]\in M$. Suppose that the following hypotheses, denoted $H_{M}([x],[z])$, hold:
  • (i)
    There exists a minimal geodesic $\gamma^{r}$ joining $[x]$ to some $q]\in M$ such that $[z]$ lies in the interior of $\gamma^{r}$.
  • (ii)
    There exists a neighborhood $U$ of $[x]$ such that for every $p]\in U$, there is a minimal geodesic joining $[z]$ to $[p]$.
  • (iii)
    For any sequence $x_{i}]\to[x]$ and any sequence $\sigma_{i}$ of minimal geodesics joining $[z]$ to $x_{i}]$, it holds that $\sigma_{i}\rightarrow\gamma$ pointwise in $M$ and $\sigma_{i}^{\prime}(0)\to\gamma^{\prime}(0)$ in $T_{[z]}M$, where $\gamma$ is the geodesic joining $[z]$ to $[x]$ given by the reverse parameterization of $\gamma^{r}|_{[x]\to[z]}$.
Then, $\gamma$ is a minimal geodesic joining $[z]$ to $[x]$ and it remains minimizing shortly beyond $[x]$.
Proof.
See Section A.2. ∎
Lemma 25.
Suppose $G\leq\operatorname{O}(d)$ is compact. For any $x,z\in P(G)$, the Riemannian manifold $[P(G)]$ satisfies the hypotheses $H_{[P(G)]}([x],[z])$ given in Lemma 24, provided $z\in Q_{x}$.
Proof.
Since $z\in Q_{x}$, there exists $g\in G$ such that $z\in g\cdot V_{x}=V_{gx}$. By Lemma 23 and its proof, the first hypothesis holds with $\gamma^{r}|_{[x]\to[z]}$ given by the image of $[gx,z]$ under the orbit map. The second hypothesis follows from (b) which entails that every pair of orbits in $[P(G)]$ are joined by some minimal geodesic in $[P(G)]$.
For the third hypothesis, consider any sequence $x_{i}]\to[x]$ and any sequence $\sigma_{i}$ of minimal geodesics joining $[z]$ to $x_{i}]$. By (b) and for each $i$, choose $g_{i}\in G$ so that $g_{i}x_{i}\in\arg\min_{q\in[x_{i}]}\|q-z\|$ and $\sigma_{i}$ are images of $z,g_{i}x_{i}]$ under the orbit map. We claim that $g_{i}x_{i}\to gx$. Once this is established, the straight lines $z,g_{i}x_{i}]$ and their derivatives at $z$ converge pointwise to the straight line $[z,gx]$ and its derivative at $z$, respectively; then these convergences descend, under the orbit map, to the desired limits $\sigma_{i}\to\gamma$ and $\sigma_{i}^{\prime}(0)\to\gamma^{\prime}(0)$.
It only remains to prove the claim $g_{i}x_{i}\to gx$. Since the sequence is bounded, it suffices to show that every convergent subsequence converges to $gx$. We fix such a convergent subsequence with indices $m_{i}$. Since $x_{m_{i}}]\to[x]$, we have $g_{m_{i}}x_{m_{i}}\to p$ for some $p\in[x]$, and we need to show that $p=gx$. By the definition of $g_{i}$, we have that $\|g_{i}x_{i}-z\|\leq\|hg_{i}x_{i}-z\|$ for each $i$ and each $h\in G$. Taking limits in $m_{i}$, we obtain that $\|p-z\|\leq\|hp-z\|$ for each $h\in G$. Thus, $p\in\arg\min_{q\in[x]}\|q-z\|=\{gx\}$, as desired. ∎
We are now ready to give an almost immediate proof of Theorem 10.
Proof of Theorem 10.
[…]
2.6 Geometric Characterization of Regular Orbits
This section is dedicated to proving Theorem 12. We proceed with similar arguments as in the previous section. To enhance readability and help absorb the core ideas, we recommend skipping the proofs of Lemmas 26, 27, 28 and 29 on the first read.
The first challenge we encounter is the non-manifold nature of $[R(G)]$. To address this issue, we desingularize the space and localize the analysis to an open neighborhood of a Voronoi cell. The following lemma could be seen as an analogue of Proposition 22. Since the proof is long and technical, we postpone it to the appendix.
Lemma 26.
Suppose $G\leq\operatorname{O}(d)$ is compact and fix $x\in R(G)$. There exists an embedded submanifold $S$ of $G$ passing through the identity $e\in G$ such that each of the following statements holds:
  1. (a)
    $S=S^{-1}$ and $S\cap G_{x}=\{e\}$.
  2. (b)
    For distinct $p\neq q\in V_{x}$, $S\cdot p\cap S\cdot q=\varnothing$.
  3. (c)
    $S\cdot V_{x}$ is an open subset of $\mathbb{R}^{d}$.
  4. (d)
    Let $M_{x}$ denote the space of equivalence classes $\{S\cdot p\}_{p\in V_{x}}$ equipped with the quotient topology. Then there exists a smooth structure and a Riemannian metric on $M_{x}$ such that the quotient map $\pi\colon S\cdot V_{x}\to M_{x}$ is a smooth Riemannian submersion.
  5. (e)
    For each $p\in V_{x}$, $S\cdot p$ is an open subset of $G\cdot p$.
  6. (f)
    For each $p\in V_{x}$, there exists an open neighborhood $W_{p}$ of $x$ in $V_{x}$ such that $\pi(W_{p})$ is open in $M_{x}$ and for $q\in W_{p}$, the minimal geodesics joining $\pi(p)$ to $\pi(q)$ in $M_{x}$ exist and are precisely the $\pi$-images of straight line distance minimizers from $p$ to $\overline{S}\cdot q$, all of which lie in $S\cdot V_{x}$.
Proof.
See Section A.3. We encourage the reader to glance at the figures therein to get a visual idea of what $S\cdot V_{x}$ and $M_{x}$ look like. ∎
This allows us to prove an analogue of Lemma 25.
Lemma 27.
Suppose $G\leq\operatorname{O}(d)$ is compact and fix $x\in R(G)$. Take $S$, $M_{x}$ and $\pi$ as in Lemma 26. The hypotheses $H_{M_{x}}(\pi(x),\pi(z))$ of Lemma 24 hold for each $z\in V_{x}$.
Proof.
By the openness of $V_{x}$ and its star convexity at $x$, as shown in 17(d), there exists $q\in V_{x}$ such that $z,x]\subseteq(q,x]\subseteq V_{x}$. Then by definition of $V_{x}$, $[q,x]$ is the unique distance minimizing straight line joining $q$ to $\overline{S}x$, and it contains the unique distance minimizing straight line $[z,x]$ joining $z$ to $\overline{S}x$. By 26(f) and since $x\in W_{q}$, we obtain that the $\pi$-image of the straight line $[q,x]$ satisfies the first hypothesis. The second hypothesis is immediate by considering the neighborhood $W_{z}$ given by 26(f). The third hypothesis follows by taking limits of distance minimizing straight line lifts $z,g_{i}x_{i}]$, where $g_{i}x_{i}\in W_{z}$, to the unique distance minimizing straight line lift $[z,x]$ as we did in the proof of Lemma 25. ∎
We also have two analogues of Lemma 23. The first concerns the (global) open Voronoi cell.
Lemma 28.
Suppose $G\leq\operatorname{O}(d)$ is compact and fix $x\in R(G)$. Take $S$, $M_{x}$ and $\pi$ as in Lemma 26. Then for each $z\in V_{x}$, there exists $q\in V_{x}$ and a minimal geodesic $\gamma$ in $M_{x}$ joining $\pi(x)$ to $\pi(q)$ such that $\pi(z)$ lies in the interior of $\gamma$.
Proof.
By the openness of $V_{x}$ and its star convexity at $x$, as shown in 17(d), there exists $q\in V_{x}$ such that $z,x]\subseteq(q,x]\subseteq V_{x}$. In particular, $\arg\min_{p\in Sx}\|q-p\|=\{x\}$ and $x\in W_{q}$, where $W_{q}$ is given by 26(f). By the definition of $W_{q}$, it follows that $\pi([q,x])$ is a minimal geodesic in $M_{x}$ joining $\pi(q)$ to $\pi(x)$ and containing $\pi(z)$ in its interior. The reverse parameterization, i.e., $\pi([x,q])$, yields the desired conclusion. ∎
Next, we have an analogue involving the local Voronoi cell.
Lemma 29.
Suppose $G\leq\operatorname{O}(d)$ is compact and fix $x\in R(G)$. Take $S$, $M_{x}$ and $\pi$ as in Lemma 26. For $z\in V_{x}$, the following statements are equivalent:
  1. (a)
    $x\in V^{loc}_{z}$.
  2. (b)
    There exists $q\in\mathbb{R}^{d}$ such that $d(q,Sz)=\|q-z\|$ and $x\in[z,q)$.
  3. (c)
    There exists $q\in V_{x}$ and a minimal geodesic $\gamma$ in $M_{x}$ joining $\pi(z)$ to $\pi(q)$ such that $\pi(x)$ lies in the interior of $\gamma$.
Proof.
[…]
$\forall q\in V,\ \ \Big|\arg\min_{p\in[z]\cap U}\|p-q\|\Big|=1.$
Since $z\in V_{x}$, we have $\{z\}=\arg\min_{p\in[z]\cap U}\|p-x\|$. By 14(b), there exists an open neighborhood $Y$ of $(x,z]$ such that $\{z\}=\arg\min_{p\in[z]\cap U}\|n-p\|$ for each $n\in Y\cap N_{z}$. By 14(c), it follows that $x\in Y$. Consequently, there exists $\varepsilon>0$ such that $q:=x+\varepsilon(z-x)\in Y\cap N_{z}$. Then $x\in[z,q)$ and $d(q,[z]\cap U)=\|q-z\|=d(q,Sz\cap U)$.
It remains to show that $d(q,Sz\cap U)=d(q,Sz)$ for small $\varepsilon$. If such $\varepsilon$ does not exist, then $d(q_{n},Sz\cap U^{c})=d(q_{n},Sz)$ for a sequence $\varepsilon_{n}\to 0$. By taking limits, we obtain that $d(x,Sz\cap U^{c})=d(x,Sz)$, which is absurd since $\arg\min_{p\in Sz}\|x-p\|=\{z\}\subseteq Sz\cap U$.
[…]
$d(z,Sq)=d(S^{-1}z,q)=d(q,Sz)=\|q-z\|.$
By 14(b), the last equality remains true when taking $q$ close enough to $x$ so that $q\in W_{z}$ and $x\in[z,q)$, where $W_{z}$ is given in 26(f). Then the $\pi$-image of the straight line $[z,q]$ minimizing distance from $z$ to $Sq$ yeilds the desired conclusion.
[…]
Then by Lemma 26 and since there is a unique minimal geodesic $\pi([z,x])$ joining $\pi(z)$ to $\pi(x)$, the horizontal lift of $\eta$ initiating from $z$ is given by a straight line $[z,uq]$ which contains $[z,x]$, where $u\in S$ and $\|uq-z\|=d(z,Sq)=d(q,Sz)$. Then $uq\in N_{x}\cap[q]=V_{x}\cap[q]$ since the straight line $z,x]\subseteq N_{x}$ is contained in $[z,uq]$. As such, $q\in V_{x}\cap V_{u^{-1}x}$ and so $u\in S\cap G_{x}=\{e\}$ by 26(a) and 17(c). The desired conclusion is satisfied by $uq=q$. ∎
We are now ready to give an almost immediate proof of Theorem 12.
Proof of Theorem 12.
[…]
$G_{x}z=R(G)\cap V_{x}\cap[z].$
By assumption, we obtain that $x\in V_{hz}^{loc}$ for each $h\in G_{x}$. By definition of $V_{hz}^{loc}$ and by 17(a), we obtain that the set $G_{x}z=\arg\min_{q\in[z]}\|q-x\|\subseteq V_{x}$ is discrete and $G_{x}$-transitive. Moreover by 18(a), we have that $G_{z}\leq G_{x}$. By transitivity and the orbit stabilizer theorem, we deduce that $G_{x}/G_{z}$ is discrete and hence finite (here, we view $G_{z}$ as the stabilizer of $z\in G_{x}z$ under the transitive action of $G_{x}$.) In particular, $\dim(G_{x})=\dim(G_{z})$. By the orbit stabilizer theorem, we get that
$\dim(G\cdot x)=\dim(G)-\dim(G_{x})=\dim(G)-\dim(G_{z})=\dim(G\cdot z).$
Since $z\in R(G)$, we obtain that $x\in R(G)$ as desired. ∎
3 Local Bilipschitzness at Regular Orbits
In this section, we provide proofs of the main results Theorems 4 and 5. The core intermediate result is given by the following lemma, whose proof is long and technical and thus postponed to the appendix.
Lemma 30.
Suppose $G\leq\operatorname{O}(d)$ is compact and define $c:=d-\max_{x\in\mathbb{R}^{d}}\dim([x])$. For $z_{1},\dots,z_{n}\in\mathbb{R}^{d}$, denote the corresponding max filter bank by $\Phi\colon\mathbb{R}^{d}/G\to\mathbb{R}^{n}$. Then the set
[…]
is semialgebraic, and it holds that
$\dim(R)\leq nd-1-\left(\left\lceil\frac{n}{\chi(G)}\right\rceil-2c+1\right).$ (4)
Proof.
See Section B.2. ∎
Here, a map $f\colon\mathbb{R}^{d}/G\to\mathbb{R}^{n}$ is said to be locally lower Lipschitz at $x\in\mathbb{R}^{d}$ if it is lower Lipschitz when restricted to a neighborhood of $x$, with respect to the quotient distance induced by $G$. The proof of Theorem 4 follows immediately from the above lemma and the fact that max filter banks are (globally) upper Lipschitz. Next, an adjustment of the proof of Theorem 5(c) in [18] gives the following proposition on the generic injectivity of max filter banks.
Proposition 31.
If $G\leq\operatorname{O}(d)$ is compact with $c:=d-\max_{x\in\mathbb{R}^{d}}\dim([x])$, then for generic $z_{1},\ldots,z_{n}\in\mathbb{R}^{d}$, the max filter bank $x]\mapsto\{\langle\hskip-2.5pt\langle[z_{i}],[x]\rangle\hskip-2.5pt\rangle\}_{i=1}^{n}$ is injective provided $n\geq 2c$.
We are now ready to give an almost immediate proof of Theorem 5.
Proof of Theorem 5.
Fix an arbitrary max filter bank $\Phi([x]):=\{\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle\}_{i=1}^{n}$. Since $\Phi$ is $\|\{z_{i}\}_{i=1}^{n}\|_{F}$-Lipschitz, it fails to be bilipschitz if and only if it fails to be lower Lipschitz. This occurs if and only if there exist sequences $x_{j}]\neq[y_{j}]$ such that
$\frac{\Phi([x_{j}])-\Phi([y_{j}])}{d([x_{j}],[y_{j}])}\rightarrow 0.$ (5)
Since (5) is symmetric and invariant under simultaneous positive dilations of $x_{j}$ and $y_{j}$, we may without loss of generality assume that $\|y_{j}\|\leq\|x_{j}\|=1$. By taking subsequences, there exist $x,y\in\mathbb{R}^{d}$ such that $x_{j}]\to[x]$ and $y_{j}]\to[y]$. Notably, since $\|x\|=1$, we have $x\in\mathbb{R}^{d}-\{0\}\subseteq R(G)$.
In the case $x]\neq[y]$, it follows that $d([x_{j}],[y_{j}])\gg 0$, and so $\Phi([x])=\Phi([y])$, meaning that $\Phi$ fails to be injective.
In the case $[x]=[y]$, it follows that $\Phi$ fails to be locally lower Lipschitz at $x\in R(G)$.
As such, if $\Phi$ is injective and locally lower Lipschitz at every $x\in R(G)$, then $\Phi$ is bilipschitz. The result now follows by combining Theorems 4 and 31. ∎
4 Discussion
In this paper, we demonstrated that sufficiently many generic templates ensure max filter banks are bilipschitz when all nonzero orbits of $G\leq\operatorname{O}(d)$ have maximal dimension, i.e., lie in $R(G)$. To achieve this, we established that max filter banks are generally locally bilipschitz on $R(G)$.
This work leaves open two key questions.
Problem 32.
  • (a)
    Is every max filter bank bilipschitz provided enough generic templates?
  • (b)
    Is every injective max filter bank bilipschitz?
Indeed, the second question is much stronger than the first. To address the first, one would need to generalize Lemma 30 to establish local lower Lipschitzness at nonregular points. A specific, unresolved example is the real irreducible representation of $\operatorname{SO}(3)$ within $\operatorname{O}(7)$.
For the stronger question, a counterexample arises if an injective max filter bank is found such that the image of $D$ in (14) is not closed. The smallest example worth investigating involves the circle group $S^{1}$ on $\mathbb{C}^{2}\times\mathbb{R}$ given by $\theta\mapsto\operatorname{diag}\{e^{i\theta},e^{i\theta},1\}$. Conversly, proving the affirmative would likely require adapting the techniques in [5] to the setting of infinite groups.
In Section 1.2, we demonstrated how max filter banks offer a theoretically desirable tool for a nearest neighbor problem in cryo-EM. Testing the hypothesis numerically could uncover potential improvements over the bispectrum embedding used in [19].
In Section 2, we derived geometric characterizations of regularity and principality using Voronoi cell decompositions. Extending these characterizations to include nonregular points remains an intriguing direction for future research.
Acknowledgments
The author thanks Dustin G. Mixon for helpful remarks, as well as Efstratios Tsoukanis for enlightening discussions.
Appendix A Riemannian geometric arguments
This section is dedicated to proving Lemmas 24 and 26. Before providing their proofs in Sections A.2 and A.3, respectively, we begin by stating necessary preliminary results in Section A.1.
A.1 Preliminaries
The first preliminary we need is for Section A.2, and it concerns subgeodesics of minimal geodesics. It is left as an exercise in Corollary 2.111 in [12]. We view the result as well-known, but we provide a proof for the sake of convenience.
Proposition 33.
Let $M$ be a Riemannian manifold and suppose that $c\colon[a,b]\to M$ is a minimal geodesic joining $c(a)$ to $c(b)$. Then $c|_{I}$ is minimal over every subinterval $I\subseteq[a,b]$.
Proof.
Let $d$ denote the geodesic distance of $M$, and let $I=[i_{0},i_{1}]\subseteq[a,b]$ be a subinterval. Let $\eta\colon I\to M$ be a unit speed piecewise $C^{1}$ curve joining $c(i_{0})$ to $c(i_{1})$ such that its length satisfies $L(\eta)\leq L(c|_{I})$. By the traingle inequality and the definition of the geodesic distance, we get that
$L(c)$ $=d(c(a),c(b))$
$\leq d(c(a),c(i_{0}))+d(c(i_{0}),c(i_{1}))+d(c(i_{1}),c(b))$
$\leq L(c|_{[a,i_{0}]})+L(\eta)+L(c|_{[i_{1},b]})$
$\leq L(c|_{[a,i_{0}]})+L(c|_{I})+L(c|_{[i_{1},b]})$
$=L(c).$
As such, $L(\eta)=L(c|_{I})$ and so $c|_{I}$ is minimal as desired. ∎
Next, we give an essential preliminary for Section A.3. It is a collection of statements regarding the orthogonal slice and tubular neighborhood geometry of compact Lie group isometric actions on manifolds. Before stating the preliminary, we define what we mean by a smooth isometric action.
Definition 34.
Let $G$ be a Lie group with identity $e$ and let $M$ be a Riemannian manifold. A smooth map $\mu\colon G\times M\to M$ is called a (left) smooth isometric action of $G$ on $M$ if:
  • (a)
    $\mu(e,x)=x$, for all $x\in M$.
  • (b)
    $\mu(g_{1},\mu(g_{2},x))=\mu(g_{1}g_{2},x)$, for all $g_{1},g_{2}\in G$ and $x\in M$.
  • (c)
    $\mu(g,\cdot)\colon M\to M$ is an isometry of $M$, for all $g\in G$.
In the following proposition, we denote $g\cdot x:=\mu(g,x)$.
Proposition 35 (Tubular Neighborhood Theorem).
Let $G$ be a compact Lie group acting smoothly and isometrically on a Riemannian manifold $M$, and fix any $x\in M$. Then $G\cdot x$ is an embedded submanifold of $M$, and there exists an open neighborhood $B$ of 0 in $(T_{x}(G\cdot x))^{\perp}$ such that each of the following statements holds:
  1. (a)
    $S_{x}:=\exp_{x}(B)$ is a $G_{x}$-invariant embedded submanifold of $M$.
  2. (b)
    $G\cdot S_{x}$ is an open neighborhood of $G\cdot x$.
  3. (c)
    Let $G\times_{G_{x}}S_{x}$ denote the orbit space of the smooth and free $G_{x}$-action on $G\times S_{x}$ given by $h\cdot(g,s):=(gh^{-1},hs)$ for $h\in G_{x}$. Then the map $\Psi_{x}\colon G\times_{G_{x}}S_{x}\to G\cdot S_{x}$, induced by $(g,s)\mapsto g\cdot s$, is a diffeomorphism.
  4. (d)
    Suppose that $H$ is an embedded submanifold of $G$ transverse to $G_{x}$ at $e$, i.e., $T_{e}H\cap T_{e}G_{x}=\{0\}$ and $T_{e}H+T_{e}G_{x}=T_{e}G$. Moreover, suppose that the multiplication map $H\times G_{x}\to H\cdot G_{x}$ is a diffeomorphism. Then the map $F\colon H\times S_{x}\to HG_{x}\times_{G_{x}}S_{x}$, induced by $(h,s)\mapsto(h,s)$, is a diffeomorphism.
For our purposes, we call any $S_{x}$ which satisfies the statements of Proposition 35 an orthogonal slice at $x$ to the left action of $G$ on $M$. This differs from the standard approach in the literature (e.g. Definition 3.47 in [1].)
Before giving a proof by references, we state a few remarks. In the proposition above, $G\times_{G_{x}}S_{x}$ is equipped with the unique smooth structure that turns the corresponding free actoin’s orbit map $G\times S_{x}\to G\times_{G_{x}}S_{x}$ into a smooth submersion. This is a special case of Theorem 3.34 in [1] which applies to general proper free actions. We also note that the name ‘Tubular Neighborhood Theorem’ stems from viewing $G\cdot S_{x}$ as an open ‘tubular’ neighborhood of $G\cdot x$.
Proof of Proposition 35.
The assertion that $G\cdot x$ is an embedded submanifold of $M$ is given by Proposition 3.41 in [1] (the result there is for so called proper actions, but we note that every smooth action of a compact Lie group is proper.) The statements of (a), (b) and (c) follow from the statement and proof of Theorem 3.57 as well as the paragraph preceeding Definition 3.72 in [1]. Lastly, (d) follows from Claim 3.52 in [1] applied to $P=G$, $F=S_{x}$, $S=H$ and $U=HG_{x}$. ∎
A.2 Proof of Lemma 24
Figure 4: This figure is an aiding illustration for Section A.2. It is an adaptation of Figure 13.2.1 in [10].
Let $d$ denote the geodesic distance of $M$, and let $\gamma^{r}:[0,d([x],[q])]\to M$ be the minimal geodesic joining $[x]$ to $[q]$, as given by the first hypothesis. By Proposition 33, its restriction $\gamma^{r}|_{[x]\to[z]}$ is a minimal geodesic joining $[x]$ to $[z]$.
Let $\gamma:[0,d([z],[x])]\to M$ be the minimal geodesic joining $[z]$ to $[x]$ defined as the reverse parameterization of $\gamma^{r}|_{[x]\to[z]}$. We aim to show that $\gamma$ remains minimizing shortly beyond $[x]$.
We begin by reframing the aim in terms of the exponential map. Since $\gamma$ is a geodesic, we have $\gamma(t)=\operatorname{exp}_{[z]}(tv)$, where $v:=\gamma^{\prime}(0)$ and $\exp\colon TM\to M$ is the Riemannian exponential map of $M$. By Proposition 5.19 in [16], the exponential map is smooth and its domain is open. It follows that $\operatorname{exp}_{[z]}(tv)$ is defined in a neighborhood of $d([z],[x])v\in T_{[z]}M$.
We seek to show that $\operatorname{exp}_{[z]}(tv)$ remains minimizing for $t$ shortly beyond $d([z],[x])$. Suppose otherwise for the sake of contradiction. We closely follow the proof of Proposition 13.2.2 in [10] which treats the case of complete manifolds. Put $t_{0}:=d([z],[x])$ and let $\{t_{0}+\varepsilon_{i}\}$ be a sequence in which $\varepsilon_{i}>0$ and $\varepsilon_{i}\to 0$. By the second hypothesis and for large $i$, there exists a sequence of minimizing geodesics $\sigma_{i}$ joining $[z]$ to $\operatorname{exp}_{[z]}((t_{0}+\varepsilon_{i})v)$, and $\sigma_{i}^{\prime}(0)\in T_{[z]}M$ is the corresponding sequence of tangent vectors at $[z]$. By the third hypothesis, $\sigma_{i}\rightarrow\gamma$ pointwise in $M$ and $\sigma_{i}^{\prime}(0)\rightarrow\gamma^{\prime}(0)$ in $T_{[z]}M$. See Figure 4 for an aiding illustration.
We show that $d\operatorname{exp}_{[z]}$ is singular at $t_{0}\gamma^{\prime}(0)$. Suppose otherwise for the sake of contradiction. Then, there exists a neighborhood $W$ of $t_{0}\gamma^{\prime}(0)$ such that $\operatorname{exp}_{[z]}\big|_{W}$ is a diffeomorphism. By definition of $\sigma_{j}$, $\gamma(t_{0}+\varepsilon_{j})=\sigma_{j}(t_{0}+\varepsilon_{j}^{\prime})$, where $\varepsilon_{j}^{\prime}\leq\varepsilon_{j}$ because $\sigma_{j}$ is minimizing. Take $\varepsilon_{j}$ sufficiently small so that $(t_{0}+\varepsilon_{j}^{\prime})\sigma_{j}^{\prime}(0)$ and $(t_{0}+\varepsilon_{j})\gamma^{\prime}(0)$ belong to $W$. Then,
$\operatorname{exp}_{[z]}((t_{0}+\varepsilon_{j})\gamma^{\prime}(0))=\operatorname{exp}_{[z]}((t_{0}+\varepsilon_{j}^{\prime})\sigma_{j}^{\prime}(0)).$
Thus $(t_{0}+\varepsilon_{j})\gamma^{\prime}(0)=(t_{0}+\varepsilon_{j}^{\prime})\sigma_{j}^{\prime}(0)$, and so $\gamma^{\prime}(0)=\sigma_{j}^{\prime}(0)$. This contradicts the assumption that $\gamma$ is no longer minimizing beyond $t_{0}$. As such, there exists nonzero $u\in T_{[z]}M$ such that $d_{t_{0}\gamma^{\prime}(0)}\operatorname{exp}_{[z]}u=0$. By Corollary 3.46 and Definition 3.72 in [12], it holds that $[z]$ and $[x]$ are conjugate along $\gamma$.
Now, let $\gamma^{r}\colon[0,d([x],[z])]\to M$ be the unique minimal geodesic joining $[x]$ to $[z]$. Since $\gamma^{r}$ is the reverse parameterization of $\gamma$, $[x]$ and $[z]$ are also conjugate along $\gamma^{r}$. However, by the first hypothesis, $\gamma^{r}$ remains minimizing shortly beyond $[z]$. This contradicts the fact that geodesics fail to be minimizing beyond conjugate points (Theorem 3.73(ii) in [12].)
A.3 Proof of Lemma 26
This section is dedicated to the proof of Lemma 26, which is both long and technical. To enhance readability and organization, the proof is divided into subsections, each corresponding to a specific part (a)-(f) of the lemma, in order. Within each subsection, we present a sequence of claims, each accompanied by proof.
Notably, the last claim in each subsection is the key result used in subsequent subsections; all other claims within the subsection serve as intermediate steps and are only relevant locally. Once the last claim in a subsection is proven, it completes the proof of the corresponding part of the lemma. To aid in visualization, we provide diagrams throughout the proof. We hope that this structure allows the reader to follow the argument linearly while minimizing the need to reference earlier claims from previous subsections.
Throughout, we view $G$ as a compact Lie subgroup and embedded submanifold of $\operatorname{O}(d)$. In particular, its Lie exponential agrees with the matrix exponential. We denote its Lie algebra by $\mathfrak{g}\subseteq\mathbb{R}^{d\times d}$ and its identity by $e:=\operatorname{id}_{\mathbb{R}^{d}}$. For a compact subgroup $H\leq G$, we denote by $H^{0}$ the connected component of $H$ which contains $e$.
A.3.1 Proof of Lemma 26(a)
We begin by constructing the desired submanifold $S$. The following is a stronger version of 26(a).
Claim A1 (Construction of $S$).
There exists an embedded submanifold $S\subseteq G$ passing through $e$ such that each of the following statements holds:
  1. (a)
    $S$ is symmetric, i.e., $S^{-1}=S$.
  2. (b)
    For each $g\in G$, $gSg^{-1}$ is transverse to $gG_{x}g^{-1}=G_{gx}$ at $e$, i.e., $T_{e}(gSg^{-1})\cap T_{e}G_{gx}=\{0\}$ and $T_{e}(gSg^{-1})+T_{e}G_{gx}=\mathfrak{g}$.
  3. (c)
    $\overline{S^{5}}\cap G_{x}\subseteq G_{x}^{0}$ and $\overline{S}\cap G_{x}=\{e\}$.
  4. (d)
    There exists an open neighborhood $T$ of 0 in $T_{e}S$ such that $\exp|_{T}\colon T\to S$ is a diffeomorphism, where $\exp\colon\mathfrak{g}\to G$ is the matrix exponential.
  5. (e)
    The multiplication map $S\times G_{x}\to S\cdot G_{x}$ is a diffeomorphism. In particular, $S\cdot G_{x}$ is an open subset of $G$.
Figure 5: This figure is an aiding illustration for the proof of Claim A1. For brevity, we define $\log(U):=\exp|_{T_{W}}^{-1}(U)$. The dashed lines connecting the left and right halves of the figure are solely for visualizing the transition by magnification from $G$ to its Lie algebra.
Proof.
See Figure 5 for an aiding illustration.
Step 1. We construct a nice neighborhood of the identity $e\in G$.
Since $G$ is a Lie group, the $5$-fold multiplication map $m_{G}\colon G\times G\times G\times G\times G\to G$ is well-defined and continuous. Since $G_{x}$ has finitely many connected components each of which is compact, there exists an open neighborhood $W$ of $G_{x}^{0}$ such that $\overline{W}\cap G_{x}=G_{x}^{0}$ and such that the matrix exponential map restricts to a diffeomorphism of an open neighborhood $T_{W}\subseteq\mathfrak{g}$ of 0 onto a subset $\exp(T_{W})\subseteq W$ that satisfies $\overline{\exp(T_{W})}\cap G_{x}=\exp(\overline{T_{W}}\cap T_{e}G_{x})$. The latter is possible since $G_{x}$ is a compact subgroup hence an embedded submanifold of $G$.
Then for a simple open neighborhood $W_{1}\times W_{2}\times W_{3}\times W_{4}\times W_{5}\subseteq m_{G}^{-1}(W)$ which contains $(e,e,e,e,e)$, define $U:=W_{1}\cap W_{2}\cap W_{3}\cap W_{4}\cap W_{5}\cap\exp(T_{W})$. One immediately observes that $U$ is a neighborhood of $e$, $\overline{U^{5}}\cap G_{x}\subseteq\overline{W}\cap G_{x}=G_{x}^{0}$, and $\overline{U}\subseteq\overline{\exp(T_{W})}$.
Step 2. We construct $S$.
Equip $G$ with a bi-invariant Riemannian metric $\beta$ so that $G_{x}$ acts freely and smoothly on $G$ by inverted right multiplication isometries. Let $(T_{e}G_{x}^{0})^{\perp}$ denote the orthogonal complement of $T_{e}G_{x}^{0}$ in $\mathfrak{g}$, with respect to $\beta$. By the Tubular Neighborhood Theorem (Proposition 35), there exists $T$, an open neighborhood of 0 in $(T_{e}G_{x}^{0})^{\perp}\cap\exp|_{T_{W}}^{-1}(U)$, such that T=T-T=T and $S:=\exp(T)\subseteq U$ is an orthogonal slice at $e$ to the aforementioned action of $G_{x}$ on $G$.
Step 3. We verify that $S$ satisfies all the properties we seek.
For (a), since T=T-T=T, it follows that $S=S^{-1}$. For (c), since $S\subseteq U$ and $\overline{U}\subseteq\overline{\exp(T_{W})}$, it follows that $\overline{S^{5}}\cap G_{x}\subseteq G_{x}^{0}$ and $\overline{S}\cap G_{x}=\{e\}$. For (d), observe that $T\subseteq T_{W}$.
Next, we prove (e). Since the aforementioned action of $G_{x}$ on $G$ is free, 35(c) entails that the right inverted (and hence noninverted) multiplication map $S\times G_{x}\to S\cdot G_{x}$ is a diffeomorphism.
Lastly, we prove (b). By 35(a), it follows that $S$ is an embedded submanifold orthogonally transverse to $G_{x}$ at $e$. Next, the adjoint map $\operatorname{Ad}_{g}\colon\mathfrak{g}\to\mathfrak{g}$, namely the derivative of conjugation $h\mapsto ghg^{-1}$ at the identity, is an isometry of $\mathfrak{g}$ with respect to the bi-invariant Riemannian metric $\beta$. As such, for each $g\in G$, it holds that $gSg^{-1}$ is orthogonally transverse to $gG_{x}g^{-1}=G_{gx}$ at $e$ as desired. ∎
A.3.2 Proof of Lemma 26(b)
Before proceeding, we need the following observation regarding stabilizers. It is an analogue of 18(c).
Claim A2.
For each $y\in R(G)$ and $p\in V_{y}$, it holds that $G_{p}\leq G_{y}$ and $G_{p}^{0}=G_{y}^{0}$. In particular, $V_{y}\subseteq R(G)$.
Proof.
The assertion $G_{p}\leq G_{y}$ follows from 18(a). In particular, $\dim(G_{p})\leq\dim(G_{y})$ and $T_{e}G_{p}\subseteq T_{e}G_{y}$. For the second assertion and since $y\in R(G)$, the orbit stabilizer theorem entails that
$\dim(G/G_{p})=\dim(G\cdot p)\leq\dim(G\cdot y)=\dim(G/G_{y}).$
As such, $\dim(G_{y})\leq\dim(G_{p})$ and so $T_{e}G_{y}=T_{e}G_{p}$. Since $G_{y}$ and $G_{p}$ are compact embedded Lie subgroups of $G$, their respective exponential maps are surjective onto their respective connected components. Thus
$G_{y}^{0}=\exp(T_{e}G_{y})=\exp(T_{e}G_{p})=G_{p}^{0}.$
The construction of $S$ and the aforementioned claim allow us to prove a stronger version of the statement of 26(b). For the following claim, see Figure 6 for an illustration.
Claim A3 (Fiber Disjointness).
Let $a$ and $b$ be nonnegative integers such that $a+b\leq 5$, and fix an arbitrary $p\in V_{x}$. For all $p_{1},p_{2}\in V_{p}$, if $p_{1}\neq p_{2}$, then $\overline{S^{a}}p_{1}\cap\overline{S^{b}}p_{2}=\varnothing$. In particular, it holds that $\overline{S^{4}}p^{\prime}\cap S\cdot V_{x}=Sp^{\prime}$, for all $p^{\prime}\in V_{x}$.
Proof.
We prove the contrapositive. If $u_{1}\in\overline{S^{a}}$ and $u_{2}\in\overline{S^{b}}$ are such that $u_{1}p_{1}=u_{2}p_{2}$, then $p_{1}=u_{1}^{-1}u_{2}p_{2}$ and so $u_{1}^{-1}u_{2}V_{p}\cap V_{p}\neq\varnothing$. By Lemmas 17 and 18(a) and A1(c), it follows that $u_{1}^{-1}u_{2}\in G_{p}\cap\overline{S^{a}}^{-1}\overline{S^{b}}\subseteq G_{x}^{0}$ (since $G_{p}\leq G_{x}$ and $\overline{S^{a}}^{-1}\cdot\overline{S^{b}}=\overline{S^{a}}\cdot\overline{S^{b}}\subseteq\overline{S^{5}}$.) By Claim A2, we get that $u_{1}^{-1}u_{2}\in G_{x}^{0}=G_{p}^{0}=G_{p_{2}}^{0}\leq G_{p_{2}}$ and so $p_{1}=u_{1}^{-1}u_{2}p_{2}=p_{2}$ as desired. ∎
Figure 6: In order from left to right, this figure is an aiding illustration for the proof of Claim A9 and the statements of Claims A5, A3 and A7, respectively.
A.3.3 Proof of Lemma 26(c)
We restate and prove 26(c) in the following claim.
Claim A4 (Open Parent Space).
$S\cdot V_{x}$ is an open subset of $\mathbb{R}^{d}$.
Proof.
Since $Q_{x}=G\cdot V_{x}$ is open by 19(d), it holds that $S\cdot V_{x}\subseteq Q_{x}$ is open in $\mathbb{R}^{d}$ if and only if $S\cdot V_{x}$ is open in $G\cdot V_{x}$. Note that $(SG_{x})\cdot V_{x}=S\cdot V_{x}$ since $G_{x}$ fixes $V_{x}$ by 17(b). Suppose for the sake of contradiction that $S\cdot V_{x}$ is not open in $G\cdot V_{x}$. Then there exist sequences $g_{n}\in G$ and $p_{n}\in V_{x}$ such that $g_{n}p_{n}\to uq\in S\cdot V_{x}$ yet $g_{n}p_{n}\notin SG_{x}\cdot V_{x}$. In particular, $g_{n}\notin SG_{x}$. By compactness of $G$, we may assume $g_{n}\to g\in G$. Note that $g\notin SG_{x}$ since $SG_{x}$ is open in $G$ by A1(e). Additionally, $p_{n}\to p:=g^{-1}uq\in V_{g^{-1}ux}\cap\overline{V}_{x}$. By 17(c), we get that $g^{-1}u\in G_{x}$ and so $g\in uG_{x}\subseteq SG_{x}$, a contradiction. ∎
A.3.4 Proof of Lemma 26(d)
By virtue of Claim A3, let $M$ denote the space of equivalence classes $\{S\cdot p\}_{p\in V_{x}}$ equipped with the quotient topology. Let $\pi\colon S\cdot V_{x}\to M$ denote the quotient map. We build up towards proving that $M$ is a topological manifold that admits a Riemannian structure that makes $\pi$ into a Riemannian submersion. We begin with describing saturations and proving that $\pi$ is an open map. For the following claim, see Figure 6 for an illustration.
Claim A5 (Open Quotient Map).
For $W\subseteq S\cdot V_{x}$, it holds that
$\pi^{-1}(\pi(W))=(S^{2}\cdot W)\cap S\cdot V_{x}.$
As a consequence, $\pi$ is an open map.
Proof.
Let $W\subseteq S\cdot V_{x}$. An element $k\in W$ has the form $k=up$ for some $u\in S$ and $p\in V_{x}$. Then, $Sp\subseteq SS^{-1}k=S^{2}k\subseteq S^{3}p$, which when combined with Claim A3 gives
$Sp\subseteq S^{2}k\cap S\cdot V_{x}\subseteq S^{3}p\cap S\cdot V_{x}=Sp.$
As such, we get that $S^{2}k\cap S\cdot V_{x}=Sp=\pi^{-1}(\pi(k))$, and so
$\pi^{-1}(\pi(W))=\cup_{k\in W}\pi^{-1}(\pi(k))=\cup_{k\in W}S^{2}k\cap S\cdot V_{x}=S^{2}\cdot W\cap S\cdot V_{x}.$
To see that $\pi$ is an open map, let $W\subseteq S\cdot V_{x}$ be open and observe that
$\pi^{-1}(\pi(W))=(\cup_{u\in S^{2}}u\cdot W)\cap S\cdot V_{x}.$
The latter set is open in the subspace topology of $S\cdot V_{x}$ since each $u\in S^{2}$ is a homeomorphism of $\mathbb{R}^{d}$. By definition of the quotient topology, we obtain that $\pi(W)$ is open as desired. ∎
As an immediate application, we prove the following claim.
Claim A6.
$M$ is Hausdorff and second countable.
Proof.
For each $p,q\in V_{x}$, we have $\overline{S}p\cap\overline{S}q=\varnothing$ by Claim A3. By compactness, these two sets are separated by open neighborhoods $O_{p}$ and $O_{q}$ in $\mathbb{R}^{d}$. Since $\pi$ is an open map, we get that $\pi(O_{p}\cap S\cdot V_{x})$ and $\pi(O_{q}\cap S\cdot V_{x})$ are open in $M$ and separate $\pi(p)$ and $\pi(q)$. This shows that $M$ is Hausdorff. Second countability is immediate from the fact that $S\cdot V_{x}$ is second countable and $\pi$ is a continuous surjective open map, as shown in Claim A5. ∎
In the aim of constructing charts on $M$, we construct nice charts in the parent space $S\cdot V_{x}$. For the following claim, see Figure 6 for an illustration.
Claim A7 (Nice Charts).
For each $u\in S$ and $p\in V_{x}$, there exists an open neighborhood $S_{up}$ of $up$ in $V_{up}\cap S\cdot V_{x}$ and an open neighborhood $T_{up}$ of 0 in $T=\exp^{-1}(S)\subseteq\mathfrak{g}$ such that each of the following holds:
  1. (a)
    $\pi|_{S_{up}}$ is injective.
  2. (b)
    $u\exp(T_{up})u^{-1}\cdot S_{up}\subseteq S\cdot V_{x}$.
  3. (c)
    The adjoint multiplication map $\Phi_{up}\colon T_{up}\times S_{up}\to u\exp(T_{up})u^{-1}\cdot S_{up}$ given by
    $\Phi_{up}(t,q):=u\exp(t)u^{-1}q$
    is a diffeomorphism. In particular, $u\exp(T_{up})u^{-1}\cdot S_{up}$ is open in $\mathbb{R}^{d}$.
  4. (d)
    The following diagram commutes,
[…]
    where $\Pi_{S_{up}}$ is given by projection onto the component of $S_{up}$.
Proof.
Step 1. We invoke the Tubular Neighborhood Theorem, i.e., Proposition 35.
In our particular case, the theorem entails that at each $up\in S\cdot V_{x}$, there exists an open $G_{up}$-invariant neighborhood $S_{up}$ of $up$ in $V_{up}\cap S\cdot V_{x}\subseteq N_{up}$ such that $G\cdot S_{up}$ is an open $G$-invariant tubular neighborhood of $G\cdot up$ and the map $\Psi_{up}\colon G\times_{G_{up}}S_{up}\to G\cdot S_{up}$, induced by multiplication $(g,s)\mapsto gs$, is a diffeomorphism.
By 35(d) and since $uSu^{-1}$ is orthogonally transverse to $G_{up}^{0}=uG_{x}^{0}u^{-1}$ and since the right multiplication map $uSu^{-1}\times G_{up}\to uS\cdot G_{p}u^{-1}$ is a diffeomorphism as shown in A1(e), the map $F\colon uSu^{-1}\times S_{up}\to uSG_{p}u^{-1}\times_{G_{up}}S_{up}$ defined by $F(r,s):=[(r,s)]$ is a diffeomorphism.
Step 2. We construct $\Phi_{up}$ via composition and we prove (b) and (c).
Recall that $S=\exp(T)$ as in A1(d), and define $\phi_{up}\colon T\times S_{up}\to G\cdot S_{up}$ by
$\phi_{up}:=\Psi_{up}\circ F\circ(\operatorname{C}_{u}\circ\exp\times\operatorname{id}_{S_{up}}),$
where $\operatorname{C}_{u}\colon G\to G$ is the diffeomorphism given by conjugation $g\mapsto ugu^{-1}$. Then $\phi_{up}$ is a diffeomorphism of $T\times S_{up}$ onto $uSu^{-1}\cdot S_{up}\subseteq\mathbb{R}^{d}$ given by $\phi_{up}(t,s)=u\exp(t)u^{-1}\cdot s$, for all $t\in T$ and $s\in S_{up}$. In particular, $uSu^{-1}\cdot S_{up}$ is open in $\mathbb{R}^{d}$ since $\dim(T\times S_{up})=\dim(uSu^{-1}\cdot S_{up})=d$.
Now, it holds that $\{0\}\times S_{up}\subseteq\phi^{-1}_{up}(uSu^{-1}\cdot S_{up}\cap S\cdot V_{x})$ and so by continuity, there exists an open neighborhood $T_{up}$ of 0 in $T$ such that $T_{up}\times S_{up}\subseteq\phi^{-1}_{up}(uSu^{-1}\cdot S_{up}\cap S\cdot V_{x})$. This proves (b). Additionally, $\Phi_{up}:=\phi_{up}|_{T_{up}\times S_{up}}$ is the diffeomorphism desired in (c).
Step 3. We prove (a).
We aim to show that $\pi|_{S_{up}}$ is injective. Let $q_{1},q_{2}\in S_{up}\subseteq S\cdot V_{x}\cap V_{up}$ and suppose that $\pi(q_{1})=\pi(q_{2})$. Then there exists $p^{\prime}\in V_{x}$ such that $q_{1},q_{2}\in Sp^{\prime}$ and so $p^{\prime}\in S^{-1}q_{j}=Sq_{j}$. Now, $u^{-1}q_{1},u^{-1}q_{2}\in V_{p}$ and $p^{\prime}\in S^{2}u^{-1}q_{1}\cap S^{2}u^{-1}q_{2}$. By Claim A3, we obtain that $u^{-1}q_{1}=u^{-1}q_{2}$ and so $q_{1}=q_{2}$ as desired.
Step 4. We prove (d).
In order to show that the diagram in (6) commutes, we show that $\pi(\Phi_{up}(t,q))=\pi(q)$ for each $t\in T_{up}$ and $q\in S_{up}\subseteq V_{up}\cap S\cdot V_{x}$. Let $p^{\prime}\in V_{x}$ be such that $q\in Sp^{\prime}$. Then $\Phi_{up}(t,q)\in S^{3}q\cap S\cdot V_{x}\subseteq S^{4}p^{\prime}\cap S\cdot V_{x}$. By Claim A3, it follows that $\Phi_{up}(t,q)\in Sp^{\prime}$ and so $\pi(\Phi_{up}(t,q))=\pi(q)=\pi(p^{\prime})$ as desired. ∎
We give a first application of these nice chart constructions.
Claim A8 (Topological Manifold).
$M$ is a topological manifold. In fact, a cover by Euclidean charts is given by the collection $\{\pi|_{S_{up}}\}_{(u,p)\in S\times V_{x}}$, where $S_{up}$ is defined in Claim A7.
Proof.
We have already shown that $M$ is Hausdorff and second countable in Claim A6. It is left show that $\{\pi|_{S_{up}}\}_{(u,p)\in S\times V_{x}}$ do indeed form a cover by Euclidean charts.
To this end, let $u\in S$ and $p\in V_{x}$ be arbitrary and put $W:=\pi(S_{up})$. Then $\pi|_{S_{up}}\colon S_{up}\to W$ is continuous and bijective by A7(a).
It is left to show that $\pi|_{S_{up}}$ is an open map. To this end, let $T_{up}$ and $\Phi_{up}$ be as defined in Claim A7, and let $Y$ be an open subset of $S_{up}$. Then by commutativity of the diagram in (6), it holds that $\pi|_{S_{up}}(Y)=\pi\circ\Phi_{up}(T_{up}\times Y)$. The latter set is open since $\pi\circ\Phi_{up}$ is an open map and $T_{up}\times Y$ is an open subset of $T_{up}\times S_{up}$. ∎
Next, we give $M$ a smooth structure so that $\pi$ becomes a smooth submersion. For the proof of the following claim, see Figure 6 for an illustration.
Claim A9 (Smooth Structure).
The charts $\{\pi|_{S_{up}}\}_{(u,p)\in S\times V_{x}}$ generate a smooth structure on $M$ which turns $\pi$ into a smooth submersion.
Proof.
Step 1. We deduce the submersion claim from smooth transitions.
Observe that the diagram in (6) is a commutative diagram of Euclidean charts that turn $\pi$ into $\Pi_{S_{up}}\colon T_{up}\times S_{up}\to S_{up}$, the projection onto the component of $S_{up}$. Due to this and once we show that the charts $\pi|_{S_{up}}$ transition smoothly, we obtain that $\pi$ is a smooth submersion as desired.
Step 2. We align two nice charts on $S\cdot V_{x}$ given in Claim A7 and setup notation.
We are left to show that the transitions between charts $\pi|_{S_{up}}$ are smooth. Let $\pi|_{S_{1}}\colon S_{1}\to W_{1}$ and $\pi|_{S_{2}}\colon S_{2}\to W_{2}$ be two arbitrary charts, where $S_{j}:=S_{u_{j}p_{j}}$ for some $u_{j}\in S$ and $p_{j}\in V_{x}$. Similarly, put $T_{j}:=T_{u_{j}p_{j}}$ and $\Phi_{j}:=\Phi_{u_{j}p_{j}}$ as given by Claim A7. Suppose that $W_{1}$ and $W_{2}$ overlap, i.e., $\pi|_{S_{1}}(q_{1})=\pi|_{S_{2}}(q_{2})\in W_{1}\cap W_{2}$ for some $q_{j}\in S_{j}$. Then there exists $u^{\prime}\in S^{2}$ such that $q_{2}=u^{\prime}q_{1}$. Now, the map $u^{\prime}\cdot\Phi_{1}\colon T_{1}\times S_{1}\to u^{\prime}\cdot u_{1}\exp(T_{1})u_{1}^{-1}\cdot S_{1}$ is a diffeomorphism, and we have that $q_{2}\in\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2})$.
Put $K_{1}:=(u^{\prime}\cdot\Phi_{1})^{-1}(\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2}))$, $K_{2}:=\Phi_{2}^{-1}(\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2}))$ and $B_{j}:=\pi|_{S_{j}}^{-1}(W_{1}\cap W_{2})$. Moreover, let $\Pi_{j}\colon T_{j}\times S_{j}\to S_{j}$ denote projection onto the component of $S_{j}$. Our main goal is to show that $\pi|_{S_{1}}^{-1}\circ\pi|_{S_{2}}$ is a diffeomorphism near $q_{2}\in B_{2}$. Then the result folows since $q_{2}\in\pi|_{S_{2}}^{-1}(W_{1}\cap W_{2})$ was chosen arbitrarily.
Step 3. We show that the following diagram is well-defined and that its bottom square commutes.
$\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2})$$W_{1}\cap W_{2}$$K_{1}$$K_{2}$$B_{2}$$B_{1}$$\scriptstyle{\pi}$$\scriptstyle{\Pi_{1}}$$\scriptstyle{u^{\prime}\cdot\Phi_{1}|_{K_{1}}}$$\scriptstyle{(u^{\prime}\cdot\Phi_{1}|_{K_{1}})^{-1}\circ\Phi_{2}|_{K_{2}}}$$\scriptstyle{\Pi_{2}}$$\scriptstyle{\Phi_{2}|_{K_{2}}}$$\scriptstyle{\pi|_{S_{1}}^{-1}\circ\pi|_{S_{2}}}$$\scriptstyle{\pi|_{S_{2}}}$$\scriptstyle{\pi|_{S_{1}}}$
In other words, we will verify that $\pi(\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2}))\subseteq W_{1}\cap W_{2}$, $\Pi_{2}(K_{2})\subseteq B_{2}$, $\Pi_{1}(K_{1})\subseteq B_{1}$ and that the bottom square commutes. We begin with showing that the outermost square commutes, i.e., $\pi\circ u^{\prime}\cdot\Phi_{1}|_{K_{1}}=\pi|_{S_{1}}\circ\Pi_{1}|_{K_{1}}$ (in particular, this entails that $\pi(\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2}))=\pi|_{S_{1}}(\Pi_{1}(K_{1}))\subseteq W_{1}$.) By the commutativity of the diagram in (6), this is equivalent to $\pi\circ u^{\prime}\cdot\Phi_{1}|_{K_{1}}=\pi\circ\Phi_{1}|_{K_{1}}$. To establish this, take any $(t,q)\in K_{1}\subseteq T_{1}\times S_{1}$ and let $p^{\prime}\in V_{x}$ be such that $\Phi_{1}|_{K_{1}}(t,q)\in Sp^{\prime}$. Then $u^{\prime}\cdot\Phi_{1}|_{K_{1}}(t,q)\in S^{3}p^{\prime}$ and so Claim A3 entails that $\pi\circ u^{\prime}\cdot\Phi_{1}|_{K_{1}}(t,q)=\pi\circ\Phi_{1}|_{K_{1}}(t,q)=\pi(p^{\prime})$ as desired.
Next, note that the middle square commutes by (6), i.e., $\pi|_{S_{2}}\circ\Pi_{2}|_{K_{2}}=\pi\circ\Phi_{2}|_{K_{2}}$. In particular, this entails that $\pi(\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2}))=\pi|_{S_{2}}(\Pi_{2}(K_{2}))\subseteq W_{2}$. As such,
$\pi(\operatorname{im}(u^{\prime}\cdot\Phi_{1})\cap\operatorname{im}(\Phi_{2}))=\pi|_{S_{1}}(\Pi_{1}(K_{1}))=\pi|_{S_{2}}(\Pi_{2}(K_{2}))\subseteq W_{1}\cap W_{2},$
and so $\Pi_{j}(K_{j})\subseteq B_{j}$ for each $j\in\{1,2\}$ as we wished to verify. In particular, the bottom square is well-defined and we are now ready to show that it commutes by the following chain of equalities:
$\pi|_{S_{1}}^{-1}\circ\pi|_{S_{2}}\circ\Pi_{2}|_{K_{2}}=\pi|_{S_{1}}^{-1}\circ\pi\circ\Phi_{2}|_{K_{2}}=\Pi_{1}|_{K_{1}}\circ(u^{\prime}\cdot\Phi_{1}|_{K_{1}})^{-1}\circ\Phi_{2}|_{K_{2}},$ (7)
where the first and second equalities hold since the middle and the outermost squares commute, respectively.
Step 4. We finish the proof by showing that $\pi|_{S_{1}}^{-1}\circ\pi|_{S_{2}}$ is a diffeomorphism near $q_{2}$.
[…]
Next, we complete the proof for the statement of 26(d).
Claim A10 (Riemannian Structure).
Equip $M$ with the symmetric bilinear form $\gamma$ defined by
$\gamma_{\pi(up)}(d\pi(up)|_{uX_{p}},d\pi(up)|_{uY_{p}}):=\langle uX_{p},uY_{p}\rangle=\langle X_{p},Y_{p}\rangle,$ (8)
for $X_{p},Y_{p}\in N_{p}$, $p\in V_{x}$ and $u\in S$. Then $\gamma$ is a well-defined (independent of $u$) smooth Riemannian metric and it turns $\pi\colon(S\cdot V_{x},\langle{,}\rangle)\to(M,\gamma)$ into a smooth Riemannian submersion whose horizontal subspaces are given by
$H_{up}:=\ker(d\pi(up))^{\perp}=N_{up},$
for $u\in S$ and $p\in V_{x}$.
Proof.
Fix arbitrary $u\in S$ and $p\in V_{x}$, and let $\pi|_{S_{up}}$ and $\Phi_{up}$ be charts as defined in Claim A7.
Step 1. We describe an orthogonal decomposition of the tangent space $T_{up}(S\cdot V_{x})$.
The linear map $\mathcal{T}_{up}\colon\mathfrak{g}\to\mathbb{R}^{d}$ given by $\mathcal{T}_{up}(\omega)=u\cdot\omega\cdot p$ has kernel $T_{e}G_{x}^{0}$ and image $\mathfrak{g}\cdot up$. By the first isomorphism theorem and since $S$ intersects $G_{x}^{0}$ transversally at $e$, we get that $\mathcal{T}_{up}|_{T_{e}S}\colon T_{e}S\to\mathfrak{g}\cdot up$ is a vector space isomorphism. As such, an arbitrary tangent vector to $up$ has an orthogonal decomposition $\mathcal{T}_{up}(\omega)\oplus uX_{p}\in(\mathfrak{g}\cdot up)\oplus N_{up}$ for some unique $X_{p}\in N_{p}$ and $\omega\in T_{e}S$.
Step 2. We show that $d\pi(up)|_{\mathcal{T}_{up}(\omega)\oplus uX_{p}}=d\pi|_{S_{up}}(up)|_{uX_{p}}$ and so $H_{up}:=\ker(d\pi(up))^{\perp}=N_{up}$.
By the commutative diagram in (6), it holds that
$d\pi(up)|_{\mathcal{T}_{up}(\omega)\oplus uX_{p}}=d\pi|_{S_{up}}\circ d\Pi_{S_{up}}\circ d\Phi_{up}^{-1}(up)|_{\mathcal{T}_{up}(\omega)\oplus uX_{p}}.$
Taking a curve $(t\omega,up+tuX_{p})\in T_{up}\times S_{up}$, we use the product rule to obtain that
$d\Phi_{up}(0,up)|_{(\omega,uX_{p})}=\left.\frac{d(u\exp(t\omega)u^{-1}(up+tuX_{p}))}{dt}\right|_{t=0}=\mathcal{T}_{up}(\omega)+uX_{p}.$
As such,
$d\Phi_{up}^{-1}(up)|_{\mathcal{T}_{up}(\omega)\oplus uX_{p}}=(\omega,uX_{p})\in T_{e}S\times N_{up}.$
Next, $d\Pi(0,up)|_{(\omega,uX_{p})}=uX_{p}\in N_{up}$. As such, $d\pi(up)|_{\mathcal{T}_{up}(\omega)\oplus uX_{p}}=d\pi|_{S_{up}}(up)|_{uX_{p}}$ and so $\ker(d\pi(up))=\mathfrak{g}\cdot up$. This proves the assertion that $\ker(d\pi(up))^{\perp}=N_{up}$.
Step 3. We prove the isometry invariance $d\pi(up)|_{uX_{p}}=d\pi(p)|_{X_{p}}$ for every $X_{p}\in N_{p}$.
Consider the curve $t\mapsto up+tuX_{p}$ where $|t|<\varepsilon$ and $\varepsilon>0$ is small enough so that $up+tuX_{p}\in S_{up}$ and $p+tX_{p}\in S_{p}$ for each $t$. Then $\pi(up+tuX_{p})=\pi(p+tX_{p})$ for each $t$. To see this, set $q:=p+tX_{p}$ and let $p^{\prime}\in V_{x}$ be such that $q\in Sp^{\prime}$. Then $uq\in S^{2}p^{\prime}\cap S\cdot V_{x}$ and so Claim A3 entails that $\pi(q)=\pi(uq)=\pi(p^{\prime})$ as desired. Taking derivatives at $t=0$, we arrive to the desired equality $d\pi(up)|_{uX_{p}}=d\pi(p)|_{X_{p}}$.
Step 4. Finishing the proof.
By Step 3, we obtain that (8) is independent of $u\in S$, i.e., $\gamma$ is well-defined. By Step 2 and (8) and since $\pi$ is a smooth submersion, we obtain that $d\pi(up)|_{H_{up}}\colon H_{up}\to T_{\pi(up)}M$ is a vector space isomorphism that pushes forward the inner product $\langle{,}\rangle$ over $H_{up}$ into $\gamma_{\pi(up)}$. This entails that $(T_{\pi(up)}M,\gamma_{\pi(up)})$ is a Hilbert space and $d\pi(up)|_{H_{up}}\colon(H_{up},\langle{,}\rangle)\to(T_{\pi(up)}M,\gamma_{\pi(up)})$ is a Hilbert space isometry. Once we show that $\gamma$ smooth, we obtain that $\gamma$ is a Riemannian metric for $M$ that turns $\pi$ into a Riemannian submersion as desired.
We are left to show that $\gamma$ is smooth. Set $n:=\dim(T_{e}S)$ and let $\omega_{1},\dots,\omega_{n}$ be a basis of $T_{e}S$. Let $p\in V_{x}$ be arbitrary. Then the diffeomorphism $\Phi_{p}\colon T_{p}\times S_{p}\to\exp(T_{p})\cdot S_{p}$ entails that $\{\omega_{i}\cdot q\}$ is a basis of $\mathfrak{g}\cdot q$ for each $q\in S_{p}$. This basis varies smoothly in $q$ and so we obtain smoothness of the map $P\colon S_{p}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ where $P_{q}x:=P(q,x)$ sends $x\in\mathbb{R}^{d}$ to its orthogonal projection onto $N_{q}$. This is due to smoothness of the Gram-Schmidt process over a smoothly varying basis. Now, put $W:=\pi(S_{p})$ and observe that $\sigma:=\pi|_{S_{p}}^{-1}\colon W\to S_{p}$ is a smooth chart that is also a local section of $\pi$, i.e., $\pi\circ\sigma=\operatorname{id}_{W}$. As such, we get that
$\gamma_{w}(v_{1},v_{2})=\gamma_{\pi(\sigma(w))}(d\pi\circ d\sigma(w)|_{v_{1}},d\pi\circ d\sigma(w)|_{v_{2}})=\langle P_{\sigma(w)}\circ d\sigma(w)|_{v_{1}},P_{\sigma(w)}\circ d\sigma(w)|_{v_{2}}\rangle,$
which proves that $\gamma_{w}(v_{1},v_{2})$ is smooth in $w\in W$ and $v_{1},v_{2}\in T_{w}M$ as desired. ∎
A.3.5 Proof of Lemma 26(e)
Recall that fibers of smooth submersions are embedded submanifolds (e.g. Corollary 5.13 in [15].) For each $p\in V_{x}$ and since $\pi$ is a smooth (Riemannian) submersion by Claim A10, the fiber $S\cdot p$ of $\pi$ is an embedded submanifold of $S\cdot V_{x}$ and hence of $\mathbb{R}^{d}$ and $[p]$ since $S\cdot V_{x}$ is an open subset of $\mathbb{R}^{d}$ as we have shown in Claim A4. Lemma 26(e) now follows by noting that $\dim(S\cdot p)=\dim(N_{p}^{\perp})=\dim([p])$.
A.3.6 Proof of Lemma 26(f)
The last two claims are geared to proving the statement of 26(f), after which we are done. For the proof of the following claim, see Figure 7 for an illustration.
Figure 7: This figure is an aiding illustration for the proof of Claim A11. While it is depicted that the curves $\eta_{n}$ converge pointwise to a limiting curve, this need not be the case.
Claim A11 (No Escape Neighborhood).
For each $p\in V_{x}$, there exists an open neighborhood $W_{p}$ of $x$ in $V_{x}$ such that $\pi(W_{p})$ is open in $M$ and for each $q\in W_{p}$, $g\in G$ and piecewise unit speed $C^{1}$ curve $\eta\colon[0,L(\eta)]\to\mathbb{R}^{d}$ joining $p$ to $gq$ with length $L(\eta)\leq d(p,\overline{S}q)=d(p,Sq)$, it holds that $\operatorname{im}(\eta)\subseteq S\cdot V_{x}$.
Proof.
We establish the second assertion. This way, the first assertion follows from shrinking $W_{p}$ so that it lies in the neighborhood $S_{x}$ over which Claim A8 applies and gives that $\pi(W_{p})$ is open in $M_{x}$. To this end, suppose for the sake of contradiction that there are sequences $g_{n}\in G$, $q_{n}\in V_{x}$ with $q_{n}\to x$ and unit speed piecewise $C^{1}$ curves $\eta_{n}\colon[0,L(\eta_{n})]\to\mathbb{R}^{d}$ joining $p$ to $g_{n}q_{n}$ with length $L(\eta_{n})\leq d(p,\overline{S}q_{n})$ such that there exists a ‘witness of escape’ $b_{n}\in\operatorname{im}(\eta_{n})\cap(S\cdot V_{x})^{c}$ for each $n$.
Since $d(p,\overline{S}q_{n})\to d(p,\overline{S}x)$ and $L(\eta_{n})\leq d(p,\overline{S}q_{n})$, we may take subsequences so that $L(\eta_{n})\rightarrow L$ and we have that
$L\leq d(p,\overline{S}x)\leq d(p,Gx)=\|p-x\|.$ (9)
We may also take subsequences so that $g_{n}\to g$. Then since we have that $\|p-g_{n}q_{n}\|\leq L(\eta_{n})$, we take limits to obtain that $\|p-gx\|\leq\|p-x\|$. Thus, $gx=x$ as $p\in V_{x}$.
Next, $\|b_{n}-p\|\leq L(\eta_{n})$ is bounded and hence we may assume $b_{n}\to b$ after taking subsequences. Then $b\in(S\cdot V_{x})^{c}$ since $b_{n}\in(S\cdot V_{x})^{c}$ and $S\cdot V_{x}$ is open by Claim A4. Since $p,x]\subseteq S\cdot V_{x}$, it follows that $b\notin[p,x]$ so that the strictness of the triangle inequality yields
$\|p-x\|<\|p-b\|+\|b-x\|-\frac{1}{M},$ (10)
for some large $M>0$. Then by (9) and (10) and for large $n$, it holds that
$L(\eta_{n})<L+\frac{1}{2M}\leq\|p-x\|+\frac{1}{2M}<\|p-b\|+\|b-x\|-\frac{1}{2M}.$
With this and since $b_{n}\in\operatorname{im}(\eta_{n})$, the following chain of inequalities holds for large $n$
$\|p-b_{n}\|+\|b_{n}-g_{n}q_{n}\|\leq L(\eta_{n})<\|p-b\|+\|b-x\|-\frac{1}{2M}.$
This contradicts the limits $b_{n}\to b$ and $g_{n}q_{n}\to gx=x$. ∎
We are alas ready to finish the proof with the following claim; see Figure 8 for a proof aid illustration.
Claim A12 (Minimal Geodesics).
Suppose that $p\in V_{x}$ and $q\in W_{p}$, where $W_{p}$ is given in Claim A11. Then, minimal geodesics from $\pi(p)$ to $\pi(q)$ in $M$ exist and are precisely the $\pi$-images of straight line distance minimizers from $p$ to $Sq$, all of which lie in $S\cdot V_{x}$.
Proof.
By Claim A11, every distance minimizing straight line $[p,uq]$ joining $p$ to $\overline{S}q$ lies in $S\cdot V_{x}$ and is as such a straight line distance minimizer from $p$ to $Sq$ since $\overline{S}q\cap S\cdot V_{x}=Sq$ by Claim A3. Since $\pi$ is a Riemannian submersion and $p-uq\in N_{uq}$ is a horizontal direction, we obtain that $\pi([p,uq])$ is a geodesic in $M$ joining $\pi(p)$ to $\pi(q)$ and of length $d(p,Sq)$ (see Proposition 2.109 in [12].)
Now, let $c_{\gamma}:[0,L(c_{\gamma})]\to M$ be a unit speed piecewise $C^{1}$ curve joining $\pi(p)$ to $\pi(q)$ where $L(c_{\gamma})$ denotes the length of $c_{\gamma}$, and suppose that $L(c_{\gamma})\leq d(p,Sq)$. The claim follows once we show that $L(c_{\gamma})=d(p,Sq)$ and that $c_{\gamma}$ is the $\pi$-image of a straight line distance minimizer from $p$ to $Sq$ in $S\cdot V_{x}$. The rest of the proof is dedicated to this objective.
Step 1. We reduce to $c_{\gamma}$ being a piecewise geodesic curve.
By covering $\operatorname{im}(c_{\gamma})$ with finitely many geodesically convex neighborhoods $\{B_{i}\}_{i=1}^{N}$ and taking a partition $0=T_{0}<T_{1}<\cdots<T_{N}=L(c_{\gamma})$ that satisfies $c_{\gamma}([T_{i-1},T_{i}])\subseteq B_{i}$ for each $i$, we may redraw each $c_{\gamma}|_{[T_{i-1},T_{i}]}$ as the unique minimal geodesic joining $c_{\gamma}(T_{i})$ to $c_{\gamma}(T_{i+1})$. If this results in $L(c_{\gamma})<d(p,Sq)$, then the mere proof of $L(c_{\gamma})=d(p,Sq)$ yields a contradiction; otherwise, the redrawing keeps $c_{\gamma}$ unaltered by the unique minimality of the pieced geodesics in each $B_{i}$. As such, we may now assume without loss of generality that $c_{\gamma}$ is a piecewise geodesic with each geodesic piece given by $c_{\gamma}|_{[T_{i-1},T_{i}]}$.
Step 2. We horizontally lift $c_{\gamma}$ piece by piece.
By local horizontal lifting of geodesics under Riemannian submersions (Proposition 2.109 in [12],) there exists a refined partition $0=t_{0}<t_{1}<\cdots<t_{n}=L(c_{\gamma})$ and for each $i\in\{0,\dots,n-1\}$, there exists $p_{i}\in\pi^{-1}(c_{\gamma}(t_{i}))\cap V_{x}$ and $q_{i+1}\in\pi^{-1}(c_{\gamma}(t_{i+1}))$ such that $p_{i},q_{i+1}]\subseteq N_{p_{i}}$ are straight lines connecting $p_{i}$ to $Sp_{i+1}$, entirely lying in $S\cdot V_{x}$, and horizontally lifting $c_{\gamma}|_{[t_{i},t_{i+1}]}$ to $S\cdot V_{x}$ starting from $p_{i}$, i.e., $\pi([p_{i},q_{i+1}])=\operatorname{im}(c_{\gamma}|_{[t_{i},t_{i+1}]})$. We take $p_{0}:=p$ and $p_{n}:=q$. In particular, observe that $\sum_{i=0}^{n-1}\|q_{i+1}-p_{i}\|=L(c_{\gamma})$.
Figure 8: This figure is an aiding illustration for the proof of Claim A12, where $n=5$. It represents the situation after we conclude that $\operatorname{im}(\eta)\subseteq S\cdot V_{x}$ in Step 4 but before we conclude that $\eta$ is a straight line and $c:=c_{\gamma}$ is a minimal geodesic in Step 5. The dotted orange lines connecting portions of $c$ in $M$ to portions of $\eta$ in $S\cdot V_{x}$ are solely for illustrating the process of horizontal lifting. The gray arrows joining $p_{i},q_{i+1}]$ to $g_{i}p_{i},g_{i+1}p_{i+1}]$ depict the notion of making minimal moves to stitch $\eta$ together.
Step 3. We take minimal moves in $G$ to attach the lifts together into a piecewise $C^{1}$ curve.
For each $i\in\{0,\dots,n-1\}$, there exists $u_{i+1}\in S$ such that $u_{i+1}p_{i+1}=q_{i+1}$. Take $u_{0}:=e$ and equip $G$ with a bi-invariant Riemannian metric $\beta$. We inductively construct a sequence of ‘moves’ $\{g_{i}\}_{i=0}^{n}\in G$. First, set $g_{0}:=e$. Assuming that we have defined $g_{i}$ for some $i\in\{0,\dots,n-1\}$, we proceed to define $g_{i+1}$.
Put $F_{i}=\{g\in G:gp_{i+1}=g_{i}q_{i+1}\}\neq\varnothing$. Then $F_{i}$ is a compact subset of $G$ and its nonemptiness is witnessed by $g_{i}u_{i+1}\in F_{i}$. With this, we define $g_{i+1}\in G$ by fixing a minimal choice
$g_{i+1}\in\arg\min_{g\in F_{i}}d_{\beta}(g,S).$ (11)
Then $g_{i+1}\in F_{i}$, i.e., $g_{i+1}p_{i+1}=g_{i}q_{i+1}$, for each $i\in\{0,\dots,n-1\}$, and the sequential concatenation of the straight line segments $g_{i}[p_{i},q_{i+1}]=[g_{i}p_{i},g_{i+1}p_{i+1}]$, for $i\in\{0,\dots,n-1\}$, yields a piecewise unit speed $C^{1}$ curve $\eta\colon[0,L(\eta)]\to\mathbb{R}^{d}$ joining $p$ to $g_{n}q$ with $L(\eta)=L(c_{\gamma})$.
Step 4. We invoke Claim A11 and show that the minimal moves are in $S$.
By construction of $\eta$, it holds that
$\|p-g_{n}q\|\leq L(\eta)=L(c_{\gamma})\leq d(p,Sq),$ (12)
and so $\operatorname{im}(\eta)\subseteq S\cdot V_{x}$ by Claim A11. In particular, $g_{i}q_{i+1}=g_{i+1}p_{i+1}\in S\cdot V_{x}$ for each $i\in\{0,\dots,n-1\}$.
From this, we shall find that $g_{i}\in S$ for each $i\in\{0,\dots,n\}$. We proceed by induction. The base case holds since $g_{0}=e\in S$. Next, suppose that $g_{i_{0}}\in S$ for some $i_{0}\in\{0,\dots,n-1\}$. Then by definition of $u_{i_{0}+1}\in S$ and since $g_{i_{0}}q_{i_{0}+1}\in S\cdot V_{x}$, we have that
$g_{i_{0}}u_{i_{0}+1}p_{i_{0}+1}=g_{i_{0}}q_{i_{0}+1}\in S\cdot V_{x}.$
Since $g_{i_{0}}u_{i_{0}+1}\in S^{2}$, it follows that $g_{i_{0}}q_{i_{0}+1}\in S^{2}p_{i_{0}+1}\cap S\cdot V_{x}$. By Claim A3, we obtain that $g_{i_{0}}q_{i_{0}+1}\in Sp_{i_{0}+1}$, and so $g_{i_{0}}q_{i_{0}+1}=up_{i_{0}+1}$ for some $u\in S$. As such, we have $u\in F_{i}\cap S$, and so $d_{\beta}(F_{i},S)=0$. By (11), we obtain that $g_{i_{0}+1}\in S$ as desired. In particular, it follows that $g_{n}\in S$.
Step 5. We finish the proof.
By (12) and since $g_{n}\in S$, we get that
$d(p,Sq)\leq\|p-g_{n}q\|\leq L(\eta)=L(c_{\gamma})\leq d(p,Sq).$
Hence, equalities hold and $\eta$ is precisely the straight line $p,g_{n}q]$, which is a minimizer of distance from $p$ to $Sq$ lying entirely in $S\cdot V_{x}$. Lastly, $\pi\circ\eta=c_{\gamma}$ since each piece of $\eta$ is given by $g_{i}[p_{i},q_{i+1}]\subseteq S\cdot V_{x}$ where $g_{i}\in S$ and $\pi(g_{i}[p_{i},q_{i+1}])=\pi([p_{i},q_{i+1}])=c_{\gamma}|_{[t_{i},t_{i+1}]}$ (the first equality follows from Claim A3 and the second equality follows by definition of $p_{i}$ and $q_{i+1}$.) ∎
Appendix B Semialgebraic geometric arguments
This section is dedicated to proving Lemma 30. We begin with a preliminary introduction to semialgebraic geometry in Section B.1. Notably, the statement of conservation of dimension (38(c)) is a key result and will be frequently referenced in Section B.2, where we provide the proof of Lemma 30.
B.1 Preliminary on Semialgebraic Sets and Groups
A basic semialgebraic set is any set of the form $\{x\in\mathbb{R}^{n}:p(x)\geq 0\}$, where $p\colon\mathbb{R}^{n}\to\mathbb{R}$ is a polynomial function. A semialgebraic set is any set obtained from some combination of finite unions, finite intersections, and complements of basic semialgebraic sets. By Proposition 2.9.10 in [6], every semialgebraic set is a finite union of manifolds. As such, the dimension of a semialgebraic set is defined as the maximum dimension of said manifolds. We say a subgroup of $\operatorname{GL}(d)$ is a semialgebraic group if it is semialgebraic as a subset of $\mathbb{R}^{d\times d}$. We say a function $\mathbb{R}^{s}\to\mathbb{R}^{t}$ is a semialgebraic function if its graph is semialgebraic as a subset of $\mathbb{R}^{s+t}$.
A pivotal observation is that, starting with a fixed collection of finitely many semialgebraic sets, one can construct new semialgebraic sets through the application of first-order logic.
Definition 36.
A first-order formula of the language of ordered fields with parameters in $\mathbb{R}$ is a formula written with a finite number of conjunctions, disjunctions, negations, and universal or existential quantifiers on variables, starting from atomic formulas which are formulas of the kind $f(x_{1},\dots,x_{n})=0$ or $g(x_{1},\dots,x_{n})>0$, where $f$ and $g$ are polynomials with coefficients in $\mathbb{R}$. The free variables of a formula are those variables of the polynomials appearing in the formula, which are not quantified.
Proposition 37 (Proposition 2.2.4 in [6]).
Let $\phi(x_{1},\dots,x_{n})$ be a first-order formula of the language of ordered fields, with parameters in $\mathbb{R}$ and with free variables $x_{1},\dots,x_{n}$. Then $\{x\in\mathbb{R}^{n}:\phi(x)\}$ is a semialgebraic set.
The above principle allows one to reveal structure in the family of semialgebraic sets. The following proposition demonstrates aspects of this, and proofs of statements therein can be found in Appendix A of [3].
Proposition 38.
The following statements regarding semialgebraic sets and functions hold:
  1. (a)
    The family of semialgebraic sets is closed under coordinate projection, complement, finite union, finite intersection and coordinate slicing.
  2. (b)
    The family of semialgebraic functions is closed under addition, multiplication, division (when defined), composition and concatenation. Moreover, fibers and images of semialgebraic functions are semialgebraic sets.
  3. (c)
    (Conservation of Dimension) If $\pi\colon\mathbb{R}^{n+d}\mapsto\mathbb{R}^{n}$ is a coordinate projection and $A$ is a semialgebraic subset of $\mathbb{R}^{n+d}$, then
    $\dim(\pi(A))\leq\dim(A)\leq\dim(\pi(A))+\max_{x\in\pi(A)}\dim(\pi^{-1}(x)\cap A).$ (13)
As mentioned in the beginning of the section, conservation of dimension is essential to many arguments in this paper. The next proposition highlights that semialgebraicity of a group is equivalent to its compactness.
Proposition 39 (Proposition 7 in [18]).
Suppose $G\leq\operatorname{O}(d)$. The following are equivalent:
  • (a)
    $G$ is topologically closed.
  • (b)
    $G$ is algebraic.
  • (c)
    $G$ is semialgebraic.
As a first application of the above propositions, we prove that the collections of principal and regular points, as well as the components of the Voronoi decomposition, form semialgebraic sets. Additionally, we demonstrate that quotient metrics and max filtering maps are semialgebraic functions.
Proposition 40.
For any compact $G\leq\operatorname{O}(d)$, the quotient metric $d([\cdot],[\cdot])\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ and the max filtering map $\langle\hskip-2.5pt\langle[\cdot],[\cdot]\rangle\hskip-2.5pt\rangle\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ are semialgebraic functions. Moreover, the sets $P(G)$, $R(G)$, $N_{x}$, $U_{x}$, $V_{x}$ and $Q_{x}$ are semialgebraic subsets of $\mathbb{R}^{d}$, for all $x\in\mathbb{R}^{d}$.
Proof.
By Proposition 39, $G$ is a semialgebraic group. Then the graph of $d([\cdot],[\cdot])$ is expressed through first order logic in the following form
$\big\{(x,z,r)\in(\mathbb{R}^{d})^{2}\times\mathbb{R}:(\forall g\in G,r\leq\|gx-z\|)\wedge(\forall\varepsilon\in\mathbb{R},\exists g\in G,\varepsilon>0\Rightarrow r+\varepsilon>\|gx-z\|)\big\},$
to which Proposition 37 applies. A similar argument applies to the max filtering map. To establish the rest of the proposition, first observe that $P(G)$ is expressed through first order logic in the following form
$P(G)=\big\{x\in\mathbb{R}^{d}:\forall p\in\mathbb{R}^{d},(\exists g\in G,gp=p\wedge gx\neq x)\vee(\forall g\in G,gp=p\iff gx=x)\big\}.$
In other words, it is either the case that $G_{p}$ is not a subgroup $G_{x}$ or otherwise, $G_{p}=G_{x}$. Next, set $D:=\max_{x\in\mathbb{R}^{d}}\dim([x])$ and let $\omega_{1},\dots,\omega_{N}$ be a basis of $\mathfrak{g}$, the Lie algebra of $G$. Then $R(G)$ is expressed with semialgebraic conditions in the following form
$R(G)=\big\{x\in\mathbb{R}^{d}:\{\omega_{i}\cdot x\}_{i=1}^{N}\text{ contains $D$ linearly independent vectors}\big\}.$
Next, for any $x\in\mathbb{R}^{d}$, $N_{x}$ is expressed with semialgebraic conditions in the following form
$N_{x}=\big\{y\in\mathbb{R}^{d}:\langle y,\omega_{i}\cdot x\rangle=0\ \forall\,i\in\{1,\dots,N\}\big\}.$
As for the Voronoi decomposition components, we define $U:=\{(x,y)\in(\mathbb{R}^{d})^{2}:y\in U_{x}\}$ and we express it with semialgebraic conditions in the following form
$U=\big\{(x,y)\in(\mathbb{R}^{d})^{2}:\|x-y\|=d([x],[y])\wedge\forall\,q\in G\cdot x-\{x\},\|q-y\|\neq d([x],[y])\big\}.$
By 38(a) and since $U_{x}$ is a coordinate slice of $U$, it follows that $U_{x}$ is a semialgebraic subset of $\mathbb{R}^{d}$.
Next, by 19(c), $Q_{x}$ is expressed with semialgebraic conditions in the following form
$Q_{x}=\big\{y\in\mathbb{R}^{d}:\exists\epsilon>0,\forall\,y^{\prime}\in\mathbb{R}^{d},|y^{\prime}-y|<\epsilon\implies(\exists p\in G\cdot x,(p,y^{\prime})\in U)\big\}.$
Lastly, 38(a) applies to the finite intersection $V_{x}=Q_{x}\cap U_{x}$. ∎
B.2 Proof of Lemma 30
This section is dedicated to the proof of Lemma 30, which is both long and technical. To enhance readability and organization, the proof is divided into subsections, each corresponding to a specific phase of the argument. The first phase is a sequence of reduction steps which force the ‘bad’ templates $\{z_{i}\}$ to lie within a common open Voronoi cell $V_{x}$. The second phase demonstrates that such bad templates lead to a nontrivial kernel for one of finitely many linear operators formed by the templates. The third and final phase shows that such noninjectivity fails to hold for sufficiently many generic templates.
Each phase is divided into a sequence of claims, each accompanied by proof. We hope this structure enables the reader to follow the argument smoothly.
For $z_{1},\dots,z_{n}\in\mathbb{R}^{d}$ and $x,y\in\mathbb{R}^{d}$ with $x]\neq[y]$, we define
$D(x,y):=\left\{\frac{\langle\hskip-2.5pt\langle[x],[z_{i}]\rangle\hskip-2.5pt\rangle-\langle\hskip-2.5pt\langle[y],[z_{i}]\rangle\hskip-2.5pt\rangle}{d([x],[y])}\right\}_{i=1}^{n}.$ (14)
Notably, $D$ is dilation invariant and $G$-invariant, meaning that
$D(x,y)=D(rgx,rg^{\prime}y),\text{ for all $r>0$ and $g,g^{\prime}\in G$}.$
B.2.1 Reduction to templates lying within a fixed open Voronoi cell
We begin by reducing to a fixed but arbitrary witness of failure.
Claim B13.
The set $R$ given by (3) is semialgebraic and to obtain the bound (4), it suffices to show that
$\dim(R_{1}(x))\leq nd-1-\left(\left\lceil\frac{n}{\chi(G)}\right\rceil-c\right),$ (15)
for all $x\in R(G)$, where
[…]
Moreover, $R_{1}(x)$ is semialgebraic.
Proof.
Fix any $p\in R(G)$ and put $N:=N_{p}$. By 19(d), $Q_{p}=G\cdot V_{p}\subseteq G\cdot N$ is dense. Since $G\cdot N$ is closed, it follows that $G\cdot N=\mathbb{R}^{d}$ and hence $N]=\mathbb{R}^{d}/G$, i.e., $N$ meets every orbit in at least one point. Since $p\in R(G)$, observe that $\dim(N)=c$.
Next, observe that $\Phi$ fails to be locally lower Lipschitz at $x\in R(G)$ if and only if there exist sequences $x_{j}],[y_{j}]\to[x]$ such that $x_{j}]\neq[y_{j}]$ and $D(x_{j},y_{j})\to 0$. Note that these conditions are dilation invariant and $G$-invariant. Then by setting $S:=R(G)\cap\mathbb{S}^{d-1}\cap N$, we have that $R$ is a coordinate projection of the following lift
[…]
By 38(a) and 40, $S$ is semialgebraic. By Proposition 37, $L$ is semialgebraic since it may be expressed in the language of first order logic as follows
$L$ $=\big\{(\{z_{i}\}_{i=1}^{n},x)\in(\mathbb{R}^{d})^{n}\times S:\forall\varepsilon\in\mathbb{R}_{>0},\exists x_{0},y_{0}\in\mathbb{R}^{d},$
$\qquad\qquad[x_{0}]\neq[y_{0}]\wedge[x_{0}],[y_{0}]\in B_{[x]}(\varepsilon)\wedge|D(x_{0},y_{0})|<\varepsilon\big\}.$
By 38(a), it follows that $R$ is semialgebraic since it is a coordinate projection of $L$. A similar first order expression shows that $R_{1}(x)$ is semialgebraic for each $x\in R(G)$. Moreover, by 38(c), we have the bounds
$\dim(R)\leq\dim(L)\leq\dim(S)+\max_{x\in S}\dim(R_{1}(x)).$
Since $\dim(S)=c-1$ and $S\subseteq R(G)$, the claim follows. ∎
Next, we reduce to templates lying within the open Voronoi diagram $Q_{x}$.
Claim B14.
Fix arbitrary $x\in R(G)$. To obtain the bound (15), it suffices to show that
$\dim(R_{2})\leq md-1-\left(\left\lceil\frac{m}{\chi(G)}\right\rceil-c\right),$ (16)
for each $m\geq 0$, where
[…]
Moreover, $R_{2}$ is semialgebraic.
Proof.
We decompose analysis based on which $z_{i}$ lies in $Q_{x}$. For $z_{1},\dots,z_{n}\in\mathbb{R}^{d}$, define $I_{\{z_{i}\}}:=\{1_{z_{i}\in Q_{x}}\}_{i=1}^{n}\in\mathbb{\{}0,1\}^{n}$. Then by 19(b), $I$ is semialgebraic in $\{z_{i}\}_{i=1}^{n}$. We obtain a finite partition $R_{1}=\bigsqcup_{I\in\{0,1\}^{n}}R_{1}^{I}$, where
$R_{1}^{I}:=R_{1}\cap\big\{\{z_{i}\}_{i=1}^{n}\in(\mathbb{R}^{d})^{n}:I_{\{z_{i}\}}=I\big\}.$
Then each $R_{1}^{I}$ is semialgebraic, and we have that $\dim(R_{1})=\max_{I\in\{0,1\}^{n}}\dim(R_{1}^{I})$. Without loss of generality, we may assume that the maximum is achieved by $I_{*}:=\{1_{j\leq m}\}_{j=1}^{n}$ for some $m\in\{0,\dots,n\}$. Then $R_{2}$ is a coordinate projection of $R_{1}^{I_{*}}$, so it is semialgebraic by 38(a). Additionally, since $R_{1}^{I_{*}}\subseteq R_{2}\times(Q_{x}^{c})^{n-m}$, it follows that
$\dim(R_{1})=\dim(R_{1}^{I_{*}})\leq\dim(R_{2})+\dim((Q_{x}^{c})^{n-m}).$
Since $Q_{x}$ is open and dense in $\mathbb{R}^{d}$ (19(d)), it holds that
$\dim((Q_{x}^{c})^{n-m})\leq(n-m)(d-1).$
The claim follows by combining the above two inequalities and noting that
$m-n\leq\left\lceil\frac{m}{\chi(G)}\right\rceil-\left\lceil\frac{n}{\chi(G)}\right\rceil.$
Next, we reduce to the case of all templates lying in the open Voronoi cell $V_{x}$.
Claim B15.
To obtain the bound (16), it suffices to show that
$\dim(R_{3})\leq mc-1-\left(\left\lceil\frac{m}{\chi(G)}\right\rceil-c\right),$ (17)
where
[…]
Moreover, $R_{3}$ is semialgebraic.
Proof.
By 17(f) and a first order logic expression, it holds that $R_{3}$ is semialgebraic. Since $Q_{x}=G\cdot V_{x}$, we have $R_{2}=G^{m}\cdot R_{3}$. Then $R_{2}$ is a coordinate projection of $L_{3}\cap(\mathbb{R}^{d}\times R_{3})$, where
$L_{3}:=\big\{(\{z_{i}\}_{i=1}^{m},\{v_{i}\}_{i=1}^{m})\in(\mathbb{R}^{d})^{m}\times V_{x}^{m}:[z_{i}]=[v_{i}]\ \,\forall\,i\in\{1,\dots,m\}\big\}.$
By 38(c) and considering the coordinate projection $\pi_{2}\colon(\mathbb{R}^{d})^{m}\times V_{x}^{m}\to V_{x}^{m}$, it follows that
$\dim(R_{2})\leq\dim(L_{3}\cap(\mathbb{R}^{d}\times R_{3}))\leq\dim(R_{3})+\max_{\{v_{i}\}_{i=1}^{m}\in V_{x}^{m}}\dim(\pi_{2}^{-1}(\{v_{i}\}_{i=1}^{m})).$
Since $\dim(\pi_{2}^{-1}(\{v_{i}\}_{i=1}^{m}))\leq m\cdot\max_{x\in\mathbb{R}^{d}}\dim(G\cdot x)=m(d-c)$, the claim follows. ∎
B.2.2 Reduction to linear operator injectivity
The following core claim establishes that the bad templates in $R_{3}$ result in a failure of injectivity for one of finitely many linear operators determined by the templates. The proof is intricate and therefore divided into six steps for clarity. Furthermore, the proof relies on the geometric characterization of regular points provided in Theorem 12.
Claim B16.
Define
$E:=\bigcap_{\{h_{l}\}_{l=1}^{m}\in(G_{x})^{m}}\bigcap_{\begin{subarray}{c}I\subseteq\{1,\dots,m\}\\ |I|=\left\lceil\frac{m}{\chi(G)}\right\rceil\end{subarray}}\big\{\{z_{i}\}_{i=1}^{m}\in N_{x}^{m}:\text{$\{\langle h_{i}z_{i},\cdot\rangle|_{N_{x}}\}_{i\in I}$ is injective}\big\}.$ (18)
Then $R_{3}\cap E=\varnothing$.
Proof.
Step 1. Establishing a concrete goal.
Suppose that there exists $\{z_{i}\}_{i=1}^{m}\in R_{3}\cap E$. By definition of $R_{3}$, there exist sequences $x_{j}],[y_{j}]\to[x]$ such that $x_{j}]\neq[y_{j}]$ and $\lim_{j\to\infty}D(x_{j},y_{j})=0$. By $G$-invariance and since $x\in Q_{x}$, we may assume $x_{j},y_{j}\in V_{x}\subseteq N_{x}$ and $x_{j},y_{j}\to x$. Note that these assumptions remain $G_{x}$-invariant.
To obtain a contradiction, it suffices to show that there exist $I\subseteq\{1,\dots,m\}$, with $|I|\geq\left\lceil\frac{m}{\chi(G)}\right\rceil$, and $u\in N_{x}$, with $\|u\|\geq 1$, such that after taking subsequences, the following convergence holds for each $i\in I$ and some $w_{i}\in[z_{i}]_{G_{x}}$:
$\frac{\langle\hskip-2.5pt\langle[x_{j}],[z_{i}]\rangle\hskip-2.5pt\rangle-\langle\hskip-2.5pt\langle[y_{j}],[z_{i}]\rangle\hskip-2.5pt\rangle}{d([x_{j}],[y_{j}])}\to\langle w_{i},u\rangle.$ (19)
This would contradict the injectivity of $\{\langle w_{i},\cdot\rangle|_{N_{x}}\}_{i\in I}$, which is guaranteed by the definition of $E$.
Step 2. Finding $w_{i}$ as limits of nearest points.
To the end of establishing (19), define the set of nearest elements in $z_{i}]$ to $y\in\mathbb{R}^{d}$ by
$N_{i}(y):=\{w\in[z_{i}]:\langle\hskip-2.5pt\langle[z_{i}],[y]\rangle\hskip-2.5pt\rangle=\langle w,y\rangle\}.$
In particular, note that $N_{i}(x)=G_{x}z_{i}$ since $z_{i}\in V_{x}$. Now, take subsequences so that the limits $\lim_{j\to\infty}N_{i}(x_{j})\subseteq G_{x}z_{i}$ and $\lim_{j\to\infty}N_{i}(y_{j})\subseteq G_{x}z_{i}$ exist for each $i$. By the pigeonhole principle, there exists $h\in G_{x}$ and $I\subseteq\{1,\dots,m\}$ such that $|I|\geq\left\lceil\frac{m}{\chi(G)}\right\rceil$ and $\lim_{j\to\infty}N_{i}(x_{j})\cap\lim_{j\to\infty}N_{i}(hy_{j})\neq\varnothing$ for each $i\in I$. By $G_{x}$-invariance of the assumptions on the sequences $x_{j}$ and $y_{j}$, we adjust the sequence $(y_{j})_{j\in\mathbb{N}}$ into $(hy_{j})_{n\in\mathbb{N}}$ and define
$L_{i}:=\lim_{j\to\infty}N_{i}(x_{j})\cap\lim_{j\to\infty}N_{i}(y_{j})\neq\varnothing,$
for each $i\in I$. Take any $w_{i}\in L_{i}\subseteq G_{x}z_{i}$ for each $i\in I$. We establish (19) with these $w_{i}$.
Figure 9: This figure is an aiding illustration for Steps 3 and 4 in the proof of Claim B16. In this plot, we have $G_{x}z_{i_{*}}=G_{x}w_{i_{*}}=\{z,hz,h^{-1}z\}$. Note that in Step 2, we had already redefined $y_{j}$ so that there are nearest neighbors of $x_{j}$ and $y_{j}$ to $z_{i_{*}}]$ (namely, $v_{z}(x_{j})$ and $v_{z}(y_{j})$, respectively) that both lie in a small neighborhood of $z=w_{i_{*}}$. This allows for analysis to go through in Steps 3 and 4, which this figure aims to visually aid in.
Step 3. Linearized analysis of the max filter (see Figure 9).
Fix arbitrary $i_{*}\in I$ and put $z:=w_{i_{*}}$. By the definition of $L_{i_{*}}$, there exist sequences $z_{j}^{x},z_{j}^{y}\to z$, where $z^{x}_{j}\in N_{i_{*}}(x_{j})$ and $z^{y}_{j}\in N_{i_{*}}(y_{j})$ for each $j$.
Since $z\in G_{x}z_{i_{*}}\subseteq V_{x}$ and $x\in R(G)$, Theorem 12 entails that $x\in V_{z}^{loc}$. By definition of $V_{z}^{loc}$, there exists open neighborhoods $V$ of $x$ and $U$ of $z$ such that
$\forall q\in V,\ \ \Big|\arg\sup_{p\in[z]\cap U}\langle p,q\rangle\Big|=1.$
By taking subsequences, we assume that $x_{j},y_{j}\in V$ and $z^{x}_{j},z^{y}_{j}\in U$ for each $j$. Furthermore, let $v_{z}\colon V\to[z]$ be the map such that $v_{z}(q)$ is the unique element in $\arg\sup_{p\in[z]\cap U}\langle p,q\rangle$. In particular, since $z\in V_{x}$, we have $v_{z}(x)=z$.
Note that $z^{x}_{j}=v_{z}(x_{j})$, $z^{y}_{j}=v_{z}(y_{j})$, and $\|x_{j}-y_{j}\|\neq 0$ since $x_{j}]\neq[y_{j}]$ for each $j$. Thus,
$\frac{\langle\hskip-2.5pt\langle[x_{j}],[z]\rangle\hskip-2.5pt\rangle-\langle\hskip-2.5pt\langle[y_{j}],[z]\rangle\hskip-2.5pt\rangle}{d([x_{j}],[y_{j}])}=\frac{\|x_{j}-y_{j}\|}{d([x_{j}],[y_{j}])}\cdot\frac{\langle x_{j},v_{z}(x_{j})\rangle-\langle y_{j},v_{z}(y_{j})\rangle}{\|x_{j}-y_{j}\|}.$ (20)
In the following step, we work on the right-hand fraction in the product and in the step after, we bound the left-hand fraction in the product.
Step 4. Analysis of the right-hand fraction in the right-hand side of (20).
We have that
$\frac{\langle x_{j},v_{z}(x_{j})\rangle-\langle y_{j},v_{z}(y_{j})\rangle}{\|x_{j}-y_{j}\|}$ $=\left\langle\frac{x_{j}-y_{j}}{\|x_{j}-y_{j}\|},v_{z}(y_{j})\right\rangle$
$\qquad\qquad-\frac{\|v_{z}(x_{j})-v_{z}(y_{j})\|}{\|x_{j}-y_{j}\|}\left\langle x_{j},\frac{v_{z}(x_{j})-v_{z}(y_{j})}{\|v_{z}(x_{j})-v_{z}(y_{j})\|}\right\rangle,$
where we define $\frac{0}{\|0\|}:=0$. By 14(b), the map $v_{z}(\cdot)$ is smooth over $V$ and hence locally upper Lipschitz there. Then since $x\in V$, there exists $C>0$ such that after taking subsequences, the following inequality holds for each $j$:
$\frac{\|v_{z}(x_{j})-v_{z}(y_{j})\|}{\|x_{j}-y_{j}\|}<C.$
Moreover, since $[z]$ is a smooth embedded submanifold of $\mathbb{R}^{d}$ and $v_{z}(x_{j}),v_{z}(y_{j})\to v_{z}(x)=z$, we get that $\frac{v_{z}(x_{j})-v_{z}(y_{j})}{\|v_{z}(x_{j})-v_{z}(y_{j})\|}\to t\in T_{z}[z]$. Note further that $x\in N_{z}$ since $z$ is a distance minimizer of $x$ to a neighborhood of $z$ in $[z]$ (14(a)). By combining the above observations, we obtain
$\left|\frac{\|v_{z}(x_{j})-v_{z}(y_{j})\|}{\|x_{j}-y_{j}\|}\left\langle x_{j},\frac{v_{z}(x_{j})-v_{z}(y_{j})}{\|v_{z}(x_{j})-v_{z}(y_{j})\|}\right\rangle\right|\leq C\left|\left\langle x_{j},\frac{v_{z}(x_{j})-v_{z}(y_{j})}{\|v_{z}(x_{j})-v_{z}(y_{j})\|}\right\rangle\right|\to C\left|\langle x,t\rangle\right|=0.$
By taking subsequences, we have $\frac{x_{j}-y_{j}}{\|x_{j}-y_{j}\|}\to u_{0}$ for some $u_{0}\in\mathbb{R}^{d}$ with $\|u_{0}\|=1$. It follows that
$\frac{\langle x_{j},v_{z}(x_{j})\rangle-\langle y_{j},v_{z}(y_{j})\rangle}{\|x_{j}-y_{j}\|}\to\langle v_{z}(x),u_{0}\rangle=\langle z,u_{0}\rangle.$ (21)
Step 5. Bounding the left-hand fraction in the right-hand side of (20).
We claim that $1\leq\liminf_{n\to\infty}\frac{\|x_{j}-y_{j}\|}{d([x_{j}],[y_{j}])}$ is upper bounded. To this end, observe that so far, $i_{*}\in I$ has been arbitrary and $u_{0}\in N_{x}$ does not depend on $i_{*}$. By definition of $E$, it holds that the set $\{w_{i}\}_{i\in I}$ is spanning. In particular, there exists $i_{0}\in I$ such that $\langle w_{i_{0}},u_{0}\rangle\neq 0$. Now, since the max filtering map $\langle\hskip-2.5pt\langle[\cdot],[w_{i_{0}}]\rangle\hskip-2.5pt\rangle$ is $\|w_{i_{0}}\|$-Lipschitz, we obtain that
$\left|\frac{\langle\hskip-2.5pt\langle[x_{j}],[w_{i_{0}}]\rangle\hskip-2.5pt\rangle-\langle\hskip-2.5pt\langle[y_{j}],[w_{i_{0}}]\rangle\hskip-2.5pt\rangle}{d([x_{j}],[y_{j}])}\right|\leq\|w_{i_{0}}\|<\infty.$
On the other hand, by (21) and since $\langle w_{i_{0}},u_{0}\rangle\neq 0$, there exists $d>0$ such that after taking subsequences, we have
$\left|\frac{\langle x_{j},v_{w_{i_{0}}}(x_{j})\rangle-\langle y_{j},v_{w_{i_{0}}}(y_{j})\rangle}{\|x_{j}-y_{j}\|}\right|>d.$
Hence by (20), we get $1\leq\frac{\|x_{j}-y_{j}\|}{d([x_{j}],[y_{j}])}\leq\frac{\|w_{i_{0}}\|}{d}<\infty$ for all $j$, which proves the subclaim of this step.
Step 6. Finishing the proof.
Take a further subsequence so that $\frac{\|x_{j}-y_{j}\|}{d([x_{j}],[y_{j}])}\to c_{0}:=\liminf_{n\to\infty}\frac{\|x_{j}-y_{j}\|}{d([x_{j}],[y_{j}])}\in[1,\infty)$. The desired convergence (19) now follows by taking $u:=c_{0}\cdot u_{0}$ and combining equations (20) and (21). ∎
B.2.3 Linear operator injectivity analysis
The proof of Lemma 30 is finished by combining the following claim with all preceding ones.
Claim B17.
To obtain the bound (17), it suffices to show that
$\dim(N_{x}^{m}-E)\leq mc-1-\left(\left\lceil\frac{m}{\chi(G)}\right\rceil-c\right).$ (22)
Moreover, the bound above holds.
Proof.
By Claim B16, we have that $R_{3}\subseteq N_{x}^{m}-E$. As such, $\dim(R_{3})\leq\dim(N_{x}^{m}-E)$, and (17) follows from (22), which we are now left to establish.
Since $N_{x}^{m}-E$ is a finite union and since the templates $\{z_{i}\}_{i\notin I}$ are unrestricted in (18), it suffices to fix $\{h_{l}\}_{l=1}^{m}\in(G_{x})^{m}$ and $I\subseteq\{1,\dots,m\}$ with $|I|=\left\lceil\frac{m}{\chi(G)}\right\rceil$ and to show that
$\dim(E_{1})\leq|I|c-1-\left(|I|-c\right),$
where
$E_{1}:=\big\{\{z_{i}\}_{i\in I}\in N_{x}^{|I|}:\text{$\{\langle h_{i}z_{i},\cdot\rangle|_{N_{x}}\}_{i\in I}$ is not injective}\big\}.$
Since $N_{x}$ is $G_{x}$-invariant (Proposition 16) and $\{h_{i}\}_{i\in I}$ is an isometry of $\mathbb{R}^{I}$, it holds that
$\dim(E_{2})=\dim(\{h_{i}\}_{i\in I}\cdot E_{1})=\dim(E_{1}),$
where
$E_{2}=\big\{\{z_{i}\}_{i\in I}\in N_{x}^{|I|}:\text{$\{\langle z_{i},\cdot\rangle|_{N_{x}}\}_{i\in I}$ is not injective}\big\}.$
By identifying $N_{x}$ with $\mathbb{R}^{c}$, we lift to the space of singular value decompositions
$F$= $\big\{\big(\{z_{i}\}_{i\in I},U,\Sigma,W\big)\in(\mathbb{R}^{c})^{|I|}\times\mathbb{R}^{c\times(c-1)}\times D_{\geq 0}^{(c-1)\times(c-1)}\times\mathbb{R}^{|I|\times(c-1)}:$
$\qquad\qquad\qquad U^{T}U=\operatorname{Id}_{c-1}\wedge W^{T}W=\operatorname{Id}_{c-1}\wedge\{\langle z_{i},\cdot\rangle|_{\mathbb{R}^{c}}\}_{i\in I}=W\Sigma U^{T}\big\},$
where $D_{\geq 0}^{(c-1)\times(c-1)}\subseteq\mathbb{R}^{(c-1)\times(c-1)}$ is the cone of diagonal matrices with nonnegative entries. Then $F$ is semialgebraic and $E_{2}$ is the projection of $F$ onto the first component $\{z_{i}\}_{i\in I}$. Let $\pi_{\sigma}$ denote the projection onto the other three components $(U,\Sigma,W)$. Then, the fibers of $\pi_{\sigma}$ are singleton. Hence by 38(c), we have $\dim(E_{2})\leq\dim(\pi_{\sigma}(F))$ and it suffices to show that
$\dim(\pi_{\sigma}(F))=|I|c-1-\left(|I|-c\right).$
Observe that
$\pi_{\sigma}(F)=\big\{\big(U,\Sigma,W\big)\in$ $\mathbb{R}^{c\times(c-1)}\times D_{\geq 0}^{(c-1)\times(c-1)}\times\mathbb{R}^{|I|\times(c-1)}:U^{T}U=W^{T}W=\operatorname{Id}_{c-1}\big\}.$
We count dimensions. By the diagonality and orthonormality constraints, it holds that $\Sigma$ has c1c-1 degrees of freedom, $U$ has c(c1)(c1)(c1)(c2)/2c(c-1)-(c-1)-(c-1)(c-2)/2 degrees of freedom and $W$ has |I|(c1)(c1)(c1)(c2)/2|I|(c-1)-(c-1)-(c-1)(c-2)/2 degrees of freedom. The total degrees of freedom are |I|c1(|I|c)|I|c-1-(|I|-c), as desired. This completes the proof of the claim and the lemma. ∎
Appendix C Non-differentiability of bilipschitz invariants
In this section, we extend Theorem 21 in [8] to the case of compact group acting on finite-dimensional real Hilbert spaces. We show that for compact groups, diffentiability of an invariant map at a nonprincipal point forbids it from being lower Lipschitz.
Proposition 41.
Suppose that $G\leq\operatorname{O}(d)$ is compact. For any $x\in P(G)^{c}$ and $G$-invariant map $f\colon\mathbb{R}^{d}\to\mathbb{R}^{n}$ that is differentiable at $x$, the following statements hold
  • (a)
    There exists a unit vector $v\in N_{x}\cap x^{\perp}$ such that $Df(x)v=0$.
  • (b)
    The induced map $f^{\downarrow}\colon\mathbb{R}^{d}/G\to\mathbb{R}^{d}$ is not lower Lipschitz.
  • (c)
    If $x\in\mathbb{S}^{d-1}$, the restriction of the induced map $f^{\downarrow}|_{[\mathbb{S}^{d-1}]}$ is not lower Lipschitz.
Proof.
First, we address (a). Since $P(G)$ and $Q_{x}=G\cdot V_{x}$ are $G$-invariant, open and dense, we can choose $p\in P(G)\cap V_{x}$. Then $G_{p}\leq G_{x}$ by 18(a), and we may pick $g\in G_{x}\setminus G_{p}$ since $x$ is not principal. By the same argument as for Theorem 21(a) in [8], it suffices to show that $\ker(g-\operatorname{id})^{\perp}\cap N_{x}\neq\{0\}$. By Proposition 16, we have that $T_{x}$ and $N_{x}$ are $g$-invariant. As such, $g$ splits as a block matrix and its $1$-eigenspace $E:=\ker(g-\operatorname{id})=(E\cap T_{x})\oplus(E\cap N_{x})$ splits into an orthogonal direct sum. It follows that $E^{\perp}\cap N_{x}=\{0\}$ if and only if $N_{x}\subseteq E$, which is false since $p\in V_{x}\cap E^{c}\subseteq N_{x}\cap E^{c}$ (note that $p\in E^{c}$ since $g\notin G_{p}$.)
For (b) and (c), the same arguments as in the proof of Theorem 21 in [8] hold as long as we establish that $d([x+tv],[x])=|t|$ for small $t$ and where $v$ is given by (a). This holds since by 17(d), we have $x+tv\in V_{x}$ for small $t$. ∎